### v1

# Rough or crumpled: Strong coupling phases of a generalized Kardar-Parisi-Zhang surface

Abstract
We study a generalized Kardar-Parisi-Zhang (KPZ) equation [D. Jana et al, Phys. Rev. E 109, L032104 (2024)], that sets the paradigm for universality in roughening of growing nonequilibrium surfaces without any conservation laws, but with competing local and nonlocal nonlinear effects. We show that such a generalized KPZ equation in two dimensions can describe a strong coupling rough or a crumpled surface, in addition to a weak coupling phase. The conformation fluctuations of such a rough surface are given by nonuniversal exponents, with orientational long-ranged order and positional short-ranged order, whereas the crumpled phase has positional and orientational short range order. Experimental and theoretical implications of these results are discussed.
The Kardar-Parisi-Zhang (KPZ) equation Kardar et al. (1986); Krug (1997); Barabási and Stanley (1995), originally proposed as a model for growing surfaces without overhangs, provides a paradigm for nonequilibrium phase transitions. It undergoes a roughening transition between a smooth phase, statistically identical to a surface described by the linear Edwards-Wilkinson (EW) equation Edwards and Wilkinson (1982), and a perturbatively inaccessible strong coupling phase Barabási and Stanley (1995), in dimension $d>2$. Perturbative dynamic renormalization group (RG) Barabási and Stanley (1995); Täuber (2014) analysis has been successful in describing this roughening transition in an $\epsilon$-expansion, where $d=2+\epsilon$, with two dimensions (2D) being the lower critical dimension of the model. In $d\leq 2$, there exists only the strong coupling phase. A KPZ surface described by a single-valued height field $h({\bf x},t)$ measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989) displays universal scaling in the long wavelength limit, characterized by $z$ and $\chi$, the dynamic and roughness exponents Barabási and Stanley (1995). These are formally defined through the correlation function
$C(r,t)\equiv\langle[h({\bf x},t)-h({\bf 0},0)]^{2}\rangle\sim r^{2\chi}\phi({r^{z}}/{t})$ (1)
in the scaling limit, where $r=|{\bf x}|$ and $\phi$ is a dimensionless scaling function of its argument Barabási and Stanley (1995). In the smooth phase, $z=2$ and $\chi=1-d/2$ Barabási and Stanley (1995) and at the roughening transition at $d=2+\epsilon,\,\epsilon>0$, $z=2$ and $\chi=0$ Barabási and Stanley (1995); Täuber (2014); Janssen et al. (1999). Failure of RG to capture the strong coupling phase has led to the development of alternative techniques. Notable are the self-consistent mode-coupling theories (MCT) Bouchaud and Cates (1993); Bhattacharjee (1998), which have predicted a variety of results on scaling within specific different calculational approaches. For example Ref. Hwa and Frey (1991) explored the scenario with $d=1$, yielding $z=3/2$, a value known exactly due to the Galilean invariance and a Fluctuation-Dissipation-Theorem Barabási and Stanley (1995). Furthermore, Ref. Doherty et al. (1994) shows that $z$ rises from $z=3/2$ at $d=1$ to 2 around $d_{U}=3.6$, the upper critical dimension of the KPZ equation. In contrast, Ref. Tu (1994) suggests $z<2$ at any finite dimension indicating $d_{U}=\infty$. Furthermore, Refs. Bhattacharjee (1998); Bhattacharjee and Bhattacharyya (2007); Basu and Frey (2004, 2009a) used MCT to predict $\chi=(4-d)/6,\,z=(8+d)/6$ that tend to agree reasonably with the numerical results Gomes-Filho et al. (2021). See also Ref. Colaiori and Moore (2001), which gives $d_{c}=4$ and dynamic exponent $z$ in different dimensions. More recently, Ref. Canet and Moore (2007) revisited the MCT for the KPZ equation and showed that solutions for the MCT equations for the KPZ equation actually have two branches or two distinct universality classes below $d<2$. At $d\geq 2$, there is only one branch that continues to $d_{U}=4$, giving scaling exponents closed to those known from numerical studies.
The dynamics of a KPZ surface depends locally on surface fluctuations, and hence cannot model nonequilibrium surfaces with nonlocal dynamics. Nonlocal effects, however, can be important in wide-ranging systems, including biological growth process Santalla and Ferreira (2018), fast nonlocal transport Aharonov and Rothman (1996) and nonlocal stabiliztion of surfaces Krug and Meakin (1991); Nicoli et al. (2009). Recently, a generalized or “active” KPZ equation (hereafter a-KPZ equation) with nonlocal effects as a conceptual model with competing local and nonlocal nonlinear effects has been proposed Jana et al. (2024)
$\frac{\partial h}{\partial t}$ $=$ $\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}$ (2)
$+$ $\lambda_{1}\int d^{d}rQ_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta.$
Here, $Q_{ij}({\bf r})$ denotes the longitudinal projection operator, which in Fourier space is expressed as $Q_{ij}({\bf k})=k_{i}k_{j}/k^{2}$, with ${\bf k}$ being a Fourier wave vector. Thus $Q_{ij}({\bf r})$, which is the inverse Fourier transform of $Q_{ij}({\bf k})$, is nonlocal in space. Physically, the term $\lambda_{1}\int d^{d}rQ_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))$ contributes to the surface velocity normal to the base plane, $v_{p}=\partial h/\partial t$, and is nonlocal in ${\bm{\nabla}}h$. Furthermore, $\eta({\bf x},t)$ is a zero-mean Gaussian-distributed white noise with a variance $\langle\eta({\bf x},t)\eta({\bf 0},0)\rangle=2D\delta^{d}({\bf x})\delta(t).$ If $\lambda_{1}=0$, Eq. (2) reduces to the usual KPZ equation. As a consequence of the competition between the local and nonlocal nonlinearities, a-KPZ equation (2) shows macroscopic properties dramatically different from the KPZ equation. It can have a stable, sub- or super-logarithmically rough surface in 2D, characterized by nonuniversal scaling exponents. More strikingly and unlike the 2D KPZ equation, it also admits a roughening transition even in 2D, again characterized by nonuniversal exponents. At $d>2$, it shows a roughening transition different from that in the KPZ equation Jana et al. (2024). Equation (2) admits a pseudo-Galilean invariance under the transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}=t$, together with $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, which ensures that the combination $\lambda+2\lambda_{1}$ does not renormalize; see also Ref. Jana et al. (2024). This further means $\chi+z=2$ exactly so long as nonlinearities remain relevant (in the scaling sense).
In this Letter, we focus on the strong coupling phases of (2). We show that these strong coupling phases can be of two distinct kinds: (i) rough phase with positional short range order (SRO), but orientational long range order (LRO), or (ii) crumpled with both positional and orientational SRO. We extract the scaling exponents of the rough phase in $d\geq 2$, which show their nonuniversal nature: The variance $\Delta\equiv\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle\sim L^{2\chi(\gamma)}$, $10<\chi(\gamma)<$ 1 where $L$ is the linear size of the surface and $\overline{h}(t)$ is the average height at time $t$. Furthermore, the time-scale of relaxation $\tau(L)\sim L^{z(\gamma)}$. Both $\chi$ and $z$ are parametrized continuously by $\gamma\equiv\lambda_{1}/\lambda$, reflecting model parameter-dependent, nonuniversal scaling. In the crumpled phase, $\chi>1$.
We now derive the above results. Nonlinear terms preclude any exact analysis. We first use a Wilson dynamic RG framework Kardar et al. (1986); Barabási and Stanley (1995); Forster et al. (1977); Täuber (2014) and build upon the results discussed in Ref. Jana et al. (2024). Dimensional analysis allows us to define two effective dimensionless coupling constants:
$g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d},\,g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}.$ (3)
Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, and $S_{d}$ is the surface area of a $d$-dimensional sphere. Considering $\lambda>0$ without any loss of generality, we notice that from (3) above that $g_{2}$ can be both positive and negative, depending upon the sign of $\gamma$. There are no one-loop corrections to $\lambda,\lambda_{1}$. The one-loop corrections to $\nu$ and $D$ can be obtained following Refs. Barabási and Stanley (1995); Forster et al. (1977); Jana et al. (2024), which allow us to calculate the one-loop RG flow equations for $g$, $g_{1}$ and hence $g_{2}$,
$\frac{dg}{dl}$ $=$ $g[2-d-g\tilde{A}(\gamma)],$ (4)
$\frac{dg_{2}}{dl}$ $=$ $\frac{g_{2}}{2}[2-d-g\tilde{A}(\gamma)],$ (5)
where $\tilde{A}(\gamma)=\frac{9}{8}\gamma^{2}+\frac{11}{8}\gamma-\frac{1}{4}$. Considering 2D, the physically relevant dimension, we focus on the RG flow lines in the $g$-$g_{2}$ plane, with (0,0) as the only RG fixed point. Its stability properties are intriguing. It is stable (i.e., attractive) along the $g_{2}$-axis, but unstable (i.e., repulsive) along the $g$-axis. This indicates the existence of a separatrix, an invariant manifold under RG in the $g$-$g_{2}$ plane, that separates the stable phase from instability. While (0,0) is the only fixed point of both (22) and (23), it may be stable or unstable. Stability requires $dg/dl<0,dg_{2}/dl<0$, so that both $g(l),\,g_{2}(l)$ flow to zero in the long wavelength limit. Else when $dg/dl>0,dg_{2}/dl>0$, both $g(l),\,g_{2}(l)$ flow away from (0,0) giving instability. These two behaviors are demarcated by the separatrices $\tilde{A}(\gamma)=0$, giving
$g_{2}=(0.161)\sqrt{g},\;g_{2}=-1.383\sqrt{g}$ (6)
for $\lambda_{1}>0$ and $\lambda_{1}<0$, respectively, as the separatrices, in $g$-$g_{2}$ plane passing through (0,0). The RG flow diagram in 2D is shown in Fig. 1(a). The red lines in Fig. 1(a) are the separatrices given in (6): the RG flow lines in the green regions bordered by the separatrices flow to the origin, giving the “weak coupling phase”. Although (0,0) is the stable fixed point in this region, the flow to (0,0) is so slow that $\nu$ and $D$ are infinitely renormalized, giving being super- or sub-logarithmically rough phases, as explored in Ref. Jana et al. (2024). In the remaining regions, both $g$ and $g_{2}$ grow monotonically with the “RG time” $l$. As soon as $g(l),\,g_{2}(l)\sim{\cal O}(1)$, which happens for a surface of finite size, the RG flow equations are outside the validity of the perturbative expansions. This is the perturbatively inaccessible “strong coupling phase”.
We first argue that the region confined between the two separatrices [red lines in Fig. 1(a)] actually consists of two subregions with distinct scaling properties (instead of just one region with a rough phase, as originally speculated in Ref. Jana et al. (2024)). To investigate that let us first consider the bare perturbation theory for $\nu$ in 2D, which is same as the RG calculations for $\nu$ except that now we extend the integrals over wave vectors down to an infrared cutoff $q_{\text{min}}=2\pi/L$. Setting $d=2$, we get for the effective $\nu_{e}$
$\nu_{e}\approx\nu_{0}+g\nu_{0}\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}\ln(L/a).$ (7)
We thus find that for $\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}<0$ or $\Bigl{(}\frac{(g_{2})^{2}}{2g}+\frac{5g_{2}}{8\sqrt{g}}\Bigl{)}<0$ i.e. $-1.25<\gamma<0$, $\nu_{e}<\nu_{0}$. In fact, for sufficiently large $L/a$, $\nu_{e}<0$, giving divergence of $\langle({\bm{\nabla}}h)^{2}\rangle$ implying surface crumpling with positional and orientational SRO. In higher dimensions $d>2$, the one loop integrals are finite and hence the lower limit can be extended to 0, corresponding to the thermodynamic limit. Then for a range of $\gamma$, $-1.25<\gamma<0$ and low enough bare $\nu_{0}$, $\nu_{e}$ can turn negative, once again implying crumpling. We will revisit the region with crumpling below, shown in purple in Fig. 1(a), by using mode coupling methods, and come to a similar conclusion, giving us confidence about the existence of a crumpled phase. See Fig. 1(b) for a schematic phase diagram in the $\lambda-\lambda_{1}$ plane.
Figure 1: (a) RG flow diagram in the $g$-$g_{2}$ plane in 2D. Stable and unstable regions are divided by the separatrices (red lines). Stable (Unstable) regions are indicated by flow lines moving towards (away from) the origin. The arrows indicate the flow directions, delineating the stable and unstable regions. In this flow diagram (0,0) is the only fixed point that is unstable along the $g$-direction, but stable along the $g_{2}$-direction. (b) Phases in the $\lambda-\lambda_{1}$ plane. The green regions are the weak coupling logarithmically rough phase, the white regions are the algebraically rough phase, and the purple regions are the crumpled phase. In (a) and (b) the solid red line represents the locus of unstable fixed points where the roughening transition occurs. (c) Variation of $\chi$ in 2D in the phase space region where it is positive, with $\gamma$ as obtained from MCT. Central region where $\chi>1$, corresponds to a crumpled surface.
To study the unstable regions in the RG flow diagrams, we now set up an MCT calculation Bhattacharjee (1998); Basu and Frey (2009b) to extract the scaling behavior; see also Basu (2004); Basu and Bhattacharjee (2005) for applications of MCT in a slightly different context. The response $G({\bf k},\omega)$ and correlation $C({\bf k},\omega)$ functions have the following scaling form
$G({\bf k},\omega)=k^{-z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}=-i\omega+\Sigma(k,\omega),$ (8)
$C({\bf k},\omega)=k^{-d-2\chi-z}f\biggl{(}\frac{\omega}{k^{z}}\biggr{)}.$ (9)
where the self-energy $\Sigma({\bf k},\omega=0)=\Gamma k^{z}$ in the zero frequency limit. Similarly for the correlator in the Lorentzian approximation, we write
$C({\bf k},\omega)=\frac{2D\Gamma k^{-2\chi-d+z}}{\omega^{2}+\Gamma^{2}k^{2z}}.$ (10)
Our argument is that the universal amplitude ratio $\Gamma^{2}/(D\lambda^{2})$ can be written down both from the one-loop diagrammatics of $G^{-1}({\bf k},\omega)$ and $C({\bf k},\omega)$. We use $\chi+z=2$, which is an exact result in the strong coupling phases (also consistent with the absence of vertex corrections at the one-loop order). Further assuming that $G^{-1}$ and $C$ are dominated by their respective one-loop contributions,which would hold if $z<2$ and $\chi>0$. Combining contributions from all the one-loop diagrams for $G({\bf k},\omega=0)$, we obtain sup
$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{\chi\,d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (11)
Next, considering the one-loop contributions to $C({\bf k},\omega=0)$, we find
$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}\biggl{[}d(d+2)+4\gamma(d+2)+12\gamma^{2}\biggl{]}}{4d(d+2)(d+3\chi-2)}.$ (12)
Now comparing RHS of Eqs. (58) and (71), we obtain a quadratic equation of $\chi$
$A\chi^{2}+B\chi+C=0.$ (13)
Here, $A=36\gamma^{2}+(6d+30)\gamma+(3d+6),\,B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},\,C=(d-2)\gamma\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]}.$ Solving (13)
$\chi=\frac{-B\pm|\sqrt{B^{2}-4AC}|}{2A}=\chi_{\pm}.$ (14)
[…]
$\chi=\frac{-B}{A}=\frac{-18\gamma^{2}-22\gamma+4}{36\gamma^{2}+42\gamma+12}.$ (15)
On the separatrices (6), $B=0$ and $A=22.77$ corresponding to $\gamma=\gamma_{c1}=0.161$ and $A=19.69$ corresponding to $\gamma=\gamma_{c2}=-1.383$; $A$ changes from these boundary values within the unstable region between the separatrices. In particular, as $A$, depending upon $\gamma$, decreases, $\chi$ grows, eventually exceeding unity. At this point, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface. (We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in a RG sense), making our theory given by Eq. (2) to breakdown.) Setting $\chi=1$ in (15), we get $\gamma=-1.043,-0.142$. Thus in the region $(-1.043<\gamma<-0.142)$ MCT predicts existence of a crumpled phase, which is the purple region in Fig. 1(a) and is slightly different from the region defined by $\nu_{e}=0$ (see above). Due to the breakdown of our theory, we cannot make any definitive conclusion about any scaling properties in the crumpled region.
Figure 2: Conjectured “Occam’s razor” global RG flows in the $g$-$g_{2}$ plane. Different speculated fixed points characterizing the rough and crumpled phases are marked. The broken lines are speculated fixed lines in the rough phase; see text.
[…]
Having ascertained the existence of rough and crumpled phases in 2D, we now speculate on the global fixed point structure including the strong coupling fixed points, which are inaccessible to perturbative RG; see Fig. 2. The RG flow lines that flow out of the fixed point (0,0) in the unstable region of the phase space should flow to one of these fixed points. To analyze these fixed points, we are first guided by the expectation that there should be a strong coupling fixed point at $g\neq 0,g_{2}=0$, which is the expected 2D KPZ string coupling fixed point and should govern the scaling laws for a 2D KPZ surface (which is rough). This is marked by a red circle (the other red circle at the origin is another fixed point that is unstable along the $g$-direction, but stable along the $g_{2}$-direction). In the present model, this fixed point is expected to be unstable against perturbation by a finite $\gamma$ or $g_{2}$. On one side of the $g$-axis, there should be a crumpled phase, which we expect to be characterized by a stable crumpled phase fixed point, marked by a blue square in Fig. 2. On the other side of the $g$-axis, there should be a rough phase. Considering the MCT prediction of nonuniversal scaling in the rough phase, we speculate a fixed line marked by broken line in Fig. 2, corresponding to the rough phase. On the side of the $g$-axis with a crumpling fixed, we also expect to have a stable rough phase fixed line for higher negative values of $g_{2}$ (or higher negative values of $\gamma$), giving nonuniversal scaling. Finally, we generally expect that for larger noises, i.e., with larger values of $D$, for which both $g$ and $g_{2}$ grow, the surface should be less and less ordered. This line of reasoning suggests that there should be a crumpling fixed point on the $g$-axis beyond the putative 2D KPZ rough phase fixed point, as well as on the $g_{2}$ axis (marked by a green triangle), indicating eventual crumpling of the surface for high enough $g$ or $g_{2}$. These are basically an “Occam”-razor style arguments: Fig. 2 has the simplest topology that naturally reduces to the known RG flow lines for small $g,\,g_{2}$, as shown in Fig. 1(a). At the same time, it gives the putative global flow lines, allowing for a transition to presumed rough and crumpled phases; see Ref. Banerjee et al. (2019); Mukherjee and Basu (2022a, b) for similar Occam’s razor arguments in different contexts.
We now consider the model in higher dimensions: $d>2$. In $d=2+\epsilon>2$, there are additional fixed points of the RG flows, other than (0,0) (stable fixed point) in the $g$-$g_{2}$ plane. In the accessible region i.e. $\tilde{A}(\gamma)>0$, (0,0) is the only stable fixed point and the long-wavelength scaling behavior matches that of the Edwards-Wilkinson (EW) equation, characterized by $z=2,\,\chi=(2-d)/2$ Jana et al. (2024). In the perturbatively inaccessible strong coupling phase, i.e., $\tilde{A}(\gamma)<0$, from an RG analysis Ref. Jana et al. (2024) argued a roughening transition with an associated $\chi=-\epsilon\,\tilde{C}(\gamma)/{\tilde{A}(\gamma)}$; $\tilde{C}(\gamma)\equiv\gamma^{2}/2+5\gamma/8$. For any $\epsilon>0$, at the roughening transition $\chi$ can grow arbitrarily large, if $\tilde{A}(\gamma)$ becomes small enough (remaining negative). As $\chi$ exceeds unity, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface. (We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in a RG sense), making our theory given by Eq. (2) to breakdown.) Setting $\chi=1$, we get a limit on $\gamma$, up to which this roughening transition between a smooth phase and an algebraically rough phase can be observed. Beyond this threshold value of $\gamma$ there is a range of $\gamma$ bounded by another threshold such that in the entire range $\tilde{A}(\gamma)$ is small enough to make $\chi>1$, suggesting that the roughening transition in this range of $\gamma$ is actually between a smooth phase and a crumpled phase (and not a rough phase)! Overall thus, similar to our discussions for 2D, we expect the strong coupling phase to consist of a rough phase with $10<\chi<$ 1 and a crumpled phase with $\chi>1$. Our further theoretical analysis below substantiate this expectation.
We now briefly discuss the MCT predictions for $d>2$. Our RG calculations already predicted a smooth phase and a roughening transition presumably to a strong coupling phase. The latter however cannot be accessed in the RG calculations. Assuming $\chi>0$, we apply MCT to extract the scaling behaviour in the strong coupling phase. Focusing on $d=3$, we get $A=36\gamma^{2}+48\gamma+15$, $B=54\gamma^{2}+78\gamma-2.5$ and $C=16\gamma^{2}+24\gamma$, giving $\chi\equiv\chi(\gamma)$. Thus, both $\chi$ and $z=2-\chi$ are nonuniversal, as they are in 2D. Secondly, by varying $\gamma$, one can make $\chi$ bigger, eventually exceeding unity. As in 2D, $\chi>1$ should imply crumpling of the surface, for which our theory breaks down. We therefore find that the strong coupling phase in 3D is similar to 2D. Similar analysis indicates analogous behavior in the strong coupling phase for any $d>2$.
We have thus studied the strong coupling phases of a generalized KPZ equation, in which local and nonlocal nonlinear effects compete (in the scaling sense). Our results from the RG and bare perturbation theory in 2D suggest that the strong coupling region in the parameter space the strong coupling phase actually is made of two distinct phases, one an algebraically rough phase with $10<\chi<$ 1 corresponding to orientational LRO but positional SRO, and another with $\chi>1$ implying a crumpled phase having both positional and orientational SRO. These are corroborated by our one-loop MCT calculations. While the crumpled phase cannot be systematically explored by our theory, our MCT predicts that the algebraic rough phase is characterized by nonuniversal scaling exponents parametrized by $\gamma$, similar to the nonuniversal exponents in the logarithmically rough phase by the RG calculations. Similar behavior including a crumpled phase and a rough phase with nonuniversal scaling exponents are predicted by MCT in 3D. The results in Ref. Jana et al. (2024) showed that chirality has the effect of suppressing the nonlinear instabilities of the RG flow. We thus expect that chiral effects together with the nonlocality should be able to suppress the crumpling of the membrane, stabilising orientational order. We note that the nonuniversal scaling exponents are a crucial outcome of the lack of vertex renormalization in one-loop MCT. We are unable to speculate whether this result protected by any “hidden” symmetry not known at present, or a fortuitous result, or will not hold at higher order perturbation theory. Thus, it would be interesting to verify our results numerically, e.g., by using pseudo-spectral methods Basu and Frey (2009a); Basu et al. (1998), or by nonperturbative methods Canet et al. (2011a, b, 2010). Nonetheless, we expect our predictions on the existence of a crumpled phase to hold true.
Acknowledgement:- A.B. thanks the SERB, DST (India) for partial financial support through the TARE scheme [file no.: TAR/2021/000170] and Alexander von Humboldt Stiftung, Germany for partial financial support through the Research Group Linkage Programme (2024).
Appendix A Renormalization group calculations
We revisit and reanalyse the renormalization group calculations on the generalized KPZ equation (2) of the main text. We follow Jana et al. (2024). We first give the path integral over $h({\bf r},t)$ and its dynamic conjugate field $\hat{h}({\bf r},t)$ Bausch et al. (1976); Täuber (2014) that is equivalent to and constructed from Eq. (2) of the main text together with the noise variance. The generating functional corresponding to Eq. (2) of the main text is given by Bausch et al. (1976); Täuber (2014)
$\mathcal{Z}=\int\mathcal{D}\hat{h}\mathcal{D}he^{-\mathcal{S}[\hat{h},h]},$ (16)
where $\hat{h}$ is the dynamic conjugate field and $\mathcal{S}$ is the action functional:
$S=-\int_{{\bf x},t}\hat{h}D\hat{h}+\int_{{\bf x},t}\hat{h}\bigg{\{}\partial_{t}h-\nu\nabla^{2}h-\frac{\lambda}{2}(\bm{\nabla}h)^{2}-\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)\bigg{\}}.$ (17)
Using the one loop diagrams from the supplemental material of Ref. Jana et al. (2024) we obtain renormalized $D$ ($D^{<}$) and renormalized $\nu$ ($\nu^{<}$):
$D^{<}=D\Biggr{[}1+\frac{\lambda^{2}D}{\nu^{3}}k_{d}\biggl{(}\frac{1}{4}+\gamma\frac{1}{d}+\gamma^{2}\frac{3}{d(d+2)}\biggl{)}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{d^{d}q}{q^{2}}\Biggr{]},$ (18)
$\nu^{<}=\nu\Biggr{[}1+\frac{\lambda^{2}D}{\nu^{3}}k_{d}\biggl{(}\frac{2-d}{4d}+\frac{\gamma}{d(d+2)}\Bigl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\Bigl{)}+\frac{\gamma^{2}}{d}\biggl{)}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{d^{d}q}{q^{2}}\Biggr{]}.$ (19)
Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, where $S_{d}$ is the surface area of a $d$-dimensional sphere and $\gamma=\frac{\lambda_{1}}{\lambda}$ as defined in the main text. Using above equations and Setting $d=2$ in the above integrals, with $b=e^{\delta l}\approx 1+\delta l$ ($l$ is the RG time) and defining dimensionless coupling constant by $g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d}$ as in the main text, we obtain the RG flow equations for $D$ and $\nu$:
$\frac{dD}{dl}=D\Biggr{[}z-d-2\chi+g\biggl{(}\frac{1}{4}+\gamma\frac{1}{d}+\gamma^{2}\frac{3}{d(d+2)}\biggl{)}\Biggr{]},$ (20)
$\frac{d\nu}{dl}=\nu\Biggr{[}z-2+g\biggl{(}\frac{2-d}{4d}+\frac{\gamma}{d(d+2)}\Bigl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\Bigl{)}+\frac{\gamma^{2}}{d}\biggl{)}\Biggr{]}.$ (21)
Also defining another dimensionless coupling constant, $g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}$ and using above equations we can write the flow equations of $g$ and $g_{2}$:
$\frac{dg}{dl}=g\Biggr{[}2-d+g\biggl{\{}\frac{1}{4}+\frac{3}{d(d+2)}\gamma^{2}+\frac{\gamma}{d}-\frac{3}{4}\frac{2-d}{d}-\frac{3}{d}\gamma^{2}-\frac{3\gamma}{d(d+2)}\biggl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\biggl{)}\biggl{\}}\Biggr{]},$
$\frac{dg_{2}}{dl}=\frac{g_{2}}{2}\Biggr{[}2-d+g\biggl{\{}\frac{1}{4}+\frac{3}{d(d+2)}\gamma^{2}+\frac{\gamma}{d}-\frac{3}{4}\frac{2-d}{d}-\frac{3}{d}\gamma^{2}-\frac{3\gamma}{d(d+2)}\biggl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\biggl{)}\biggl{\}}\Biggr{]}.$
To the lowest order in $d-d_{c}$, where $d_{c}=2$ is the critical dimension,
$\frac{dg}{dl}$ $=$ $g[2-d-g\tilde{A}(\gamma)],$ (22)
$\frac{dg_{2}}{dl}$ $=$ $\frac{g_{2}}{2}[2-d-g\tilde{A}(\gamma)].$ (23)
Where $\tilde{A}(\gamma)=\frac{9}{8}\gamma^{2}+\frac{11}{8}\gamma-\frac{1}{4}$ as given in the main text.
Appendix B MCT Calculations
We start with the general definitions of the correlation $C({{\bf k}},\omega)$ and response functions $G({{\bf k}},\omega)$
$\delta^{(d)}({\bf k+k^{{}^{\prime}}})\delta(\omega+\omega^{{}^{\prime}})G({{\bf k}},\omega)=\biggl{\langle}\frac{\delta h({\bf k},\omega)}{\delta\eta({\bf k^{{}^{\prime}}},\omega^{{}^{\prime}})}\biggr{\rangle}$ (24)
$\delta^{(d)}({\bf k+k^{{}^{\prime}}})\delta(\omega+\omega^{{}^{\prime}})C({{\bf k}},\omega)=\langle h({\bf k},\omega)h({\bf k^{{}^{\prime}}},\omega^{{}^{\prime}})\rangle.$ (25)
The response and correlation functions are assumed to have the following scaling form
$G({\bf k},\omega)=k^{-z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}$ (26)
$C({\bf k},\omega)=k^{-d-2\chi-z}f\biggl{(}\frac{\omega}{k^{z}}\biggr{)},$ (27)
where $z$ is the dynamic exponent and $\chi$ is the roughness exponent. The Dyson’s equation for the self energy in the scaling limit is
$G^{-1}({\bf k},\omega)=-i\omega+\Sigma({\bf k},\omega).$ (28)
The zero frequency self energy or the relaxation rate has the form
$\Sigma({\bf k},0)=\Gamma k^{z}.$ (29)
Here $\Gamma$ is basically renormalised or effective $\nu$. The correlation function in Lorentzian approximation can be written as
$C({\bf k},\omega)=\frac{2D\Gamma k^{-2\chi-d+z}}{\omega^{2}+\Gamma^{2}k^{2z}}.$ (30)
Furthermore, the zero-frequency correlation function is
$C({\bf k},0)=\frac{2D}{\Gamma}k^{-2\chi-d-z}.$ (31)
Figure 3: One-loop Feynman diagrams that contribute to the zero frequency self energy.
Figure 3(a) shows the one-loop diagrammatic correction to the self energy at ${\cal O}(\lambda)^{2}$ that comes with a symmetry factor of $8$ and contributes
$\Bigl{(}\frac{\lambda}{2}\Bigl{)}^{2}\times\frac{1}{2!}\times 8\int_{{\bf q},\Omega}{\bf q\cdot(k-q)(q\cdot k)}C({\bf q},\Omega)G({\bf k-q},-\Omega)=\lambda^{2}\int_{{\bf q},\Omega}\frac{{\bf(k-q)\cdot q\times(k\cdot q)}\times 2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(i\Omega+\Gamma({\bf k-q})^{z})}$ (32)
After performing the $\Omega$ integral and making ${\bf q}\rightarrow{\bf(q+\frac{k}{2})}$ above integral becomes,
$I$ $=\frac{\lambda^{2}D}{\Gamma}\int_{\bf q}\Bigl{(}{\bf q+\frac{k}{2}}\Bigl{)}^{-2\chi-d+z}{\bf\biggl{[}\Bigl{(}\frac{k}{2}-q\Bigl{)}\cdot\Bigl{(}q+\frac{k}{2}\Bigl{)}\biggl{]}}\frac{{\bf\biggl{[}k\cdot\Bigl{(}q+\frac{k}{2}\Bigl{)}\biggl{]}}}{{\bf\Bigl{(}q+\frac{k}{2}\Bigl{)}}^{z}+{\bf\Bigl{|}q-\frac{k}{2}\Bigl{|}}^{z}}$
$=\frac{\lambda^{2}D}{2\Gamma}\frac{2\chi+d}{2}\int_{\bf q}\frac{q^{2}({\bf k\cdot q})q^{-2\chi-d}}{q^{z}}\frac{({\bf k\cdot q})}{q^{2}}-\frac{\lambda^{2}D}{2\Gamma}k^{2}\int_{\bf q}\frac{q^{2}}{2q^{z}}q^{-2\chi-d}$ (33)
After simplifications, the first contribution in the second line of the above integral becomes
$\frac{\lambda^{2}D}{2\Gamma}\times\frac{2\chi+d}{2}\times\frac{k_{d}}{d}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
Similarly, the second contribution gives
$\frac{\lambda^{2}D}{2\Gamma}\times\frac{k_{d}}{2}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
Adding the above contributions, we get the ${\cal O}(\lambda^{2})$ correction to the zero frequency self energy
$I=\frac{\lambda^{2}D}{4\Gamma}\frac{2}{d}k_{d}k^{2-\chi}.$ (34)
Figure 3(b) shows the one-loop contribution to the self energy at ${\cal O}(\lambda_{1})^{2}$ with a symmetry factor of $8$ and contributes
$I$ $=\frac{\lambda_{1}^{2}}{2!}\times 8\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}q_{i}(k-q)_{j}\frac{(q-k)_{m}(q-k)_{n}}{(q-k)^{2}}q_{m}k_{n}C({\bf q},\Omega)\times G({\bf k-q},-\Omega)$
$=\frac{4\lambda_{1}^{2}k_{i}k_{j}k_{n}}{k^{2}}\int_{{\bf q},\Omega}q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}\frac{1}{{\bf(q-k)^{2}}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})}\times\frac{1}{i\Omega+\Gamma{\bf(k-q)}^{z}}$ (35)
After performing the $\Omega$ integral, we get
$I=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}q^{-2\chi-d}}{{\bf(q-k)}^{2}(q^{z}+{\bf|k-q|}^{z})}.$ (36)
Then shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$, above integral becomes
$\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q-\frac{k}{2})_{m}(q-\frac{k}{2})_{n}(q+\frac{k}{2})_{m}}{{\bf(q-\frac{k}{2})}^{2}({\bf|q+\frac{k}{2}|}^{z}+{\bf|q-\frac{k}{2}|}^{z})}{\bf(q+\frac{k}{2})}^{-2\chi-d}.$ (37)
Next, we extract the ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from the numerator. Then above integral $I$ can be split into three parts $I_{1}$, $I_{2}$ and $I_{3}$, with
$I_{1}$ $=-\frac{4\lambda_{1}^{2}D}{2\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}q_{i}q_{j}q_{m}q_{n}q^{2-2\chi-d-z-4}=-\frac{6\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{1}{\chi}k_{d}k^{-\chi+2},$ (38)
$I_{2}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{j}q_{m}q_{n}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}=\frac{3\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{2\chi+d}{\chi}k_{d}k^{-\chi+2},$ (39)
where we have used the following identity Yakhot and Orzag (1986),
$k_{i}k_{j}k_{m}k_{n}\int d^{d}qf(q^{2})q_{i}q_{j}q_{m}q_{n}=k_{i}k_{j}k_{m}k_{n}\frac{[\delta_{ij}\delta_{mn}+\delta_{im}\delta_{jn}+\delta_{in}\delta_{jm}]}{d(d+2)}\int d^{d}qf(q^{2})q^{4},$ (40)
and
$I_{3}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}k_{i}k_{j}\int_{\bf q}\frac{q_{i}q_{j}q^{2}q^{-2\chi-d}}{4q^{2}q^{z}}=\frac{\lambda_{1}^{2}D}{\Gamma d}\frac{1}{\chi}k_{d}k^{-\chi+2}.$ (41)
Then, $I_{3}$ is evaluated by using the well-known relation Yakhot and Orzag (1986)
$k_{i}k_{j}\int d^{d}qf(q^{2})q_{i}q_{j}=k_{i}k_{j}\times\frac{[\delta_{ij}]}{d}\int d^{d}qf(q^{2})q^{2}.$ (42)
Adding $I_{1}$, $I_{2}$ and $I_{3}$ we get contribution to zero frequency self energy from vertices $\lambda_{1}$-$\lambda_{1}$
$-\frac{\lambda_{1}^{2}D}{\Gamma}\frac{k_{d}}{d}\frac{1}{\chi}k^{2-\chi}\biggl{(}\frac{4-4d-6\chi}{d+2}\biggl{)}.$ (43)
Figure 3(c) gives one of the two one-loop corrections to the self energy at ${\cal O}(\lambda_{1}\lambda)$,with a contribution
$I=\frac{4\lambda\lambda_{1}}{2!}\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}({\bf q\cdot k})\frac{1}{i\Omega+\Gamma{\bf|k-q|}^{z}}\frac{2D\Gamma q^{-2\chi-d+z}}{\Omega^{2}+\Gamma^{2}q^{2z}}.$ (44)
After performing the $\Omega$ integral and shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$, we get
$\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}\frac{(q+\frac{k}{2})^{-2\chi-d}}{{\bf(q+\frac{k}{2})}^{z}+{\bf|\frac{k}{2}-q|}^{z}}.$ (45)
$=-\frac{2\lambda\lambda_{1}D}{\Gamma}\times\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}q_{i}q_{j}\bigl{(}q_{m}+\frac{k_{m}}{2}\bigl{)}\frac{q^{-2\chi-d}}{2q^{z}}\biggl{(}1+\frac{{\bf k\cdot q}}{q^{2}}\biggl{)}^{-(\frac{2\chi+d}{2})}$ (46)
$=\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\frac{2\chi+d}{2}\int_{{\bf q}}\frac{q_{i}q_{j}q_{m}q_{n}}{2q^{2}q^{z}}q^{-2\chi-d}-\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}}{2}\int_{{\bf q}}\frac{q_{i}q_{j}}{2q^{z}}q^{-2\chi-d}$ (47)
$=\frac{\lambda\lambda_{1}D}{\Gamma}3k^{2}k_{d}\frac{2\chi+d}{2}\frac{1}{d(d+2)}\int_{{\bf q}}q^{4-2\chi-d-z-2+d-1}-\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k^{2}k_{d}}{d}\int_{{\bf q}}q^{2-2\chi-d-z+d-1}.$ (48)
In evaluation of these integrals we used identity (40) and (42). After some simplifications, the above integral gives
$-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{2d}\frac{k^{2-\chi}}{\chi}\biggl{(}\frac{2-6\chi-2d}{d+2}\biggl{)}.$ (49)
Finally, Fig. 3(d) gives the second one-loop correction to the self energy that is also ${\cal O}(\lambda\lambda_{1})$, but is distinct from the one in Fig. 3(c).
$I=\frac{4\lambda\lambda_{1}}{2!}\int_{{\bf q},\Omega}{\bf(q\cdot(k-q))}\frac{(q-k)_{i}(q-k)_{j}}{{\bf(q-k)}^{2}}q_{i}k_{j}\times\frac{1}{i\Omega+\Gamma{\bf|k-q|}^{z}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{\Omega^{2}+\Gamma^{2}q^{2z}}.$ (50)
After performing the $\Omega$ integral above integral is,
$\frac{2\lambda\lambda_{1}D}{\Gamma}K_{j}\int_{{\bf q}}\frac{{\bf(q\cdot(k-q))}(q-k)_{i}(q-k)_{j}q_{i}q^{-2\chi-d}}{{\bf(q-k)}^{2}\times(q^{z}+{\bf|k-q|}^{z})}.$ (51)
After shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$ above integral reduces to
$\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}(q+\frac{k}{2})_{m}(\frac{k}{2}-q)_{m}(q-\frac{k}{2})_{i}(q-\frac{k}{2})_{j}(q+\frac{k}{2})_{i}\times\frac{{\bf(q+\frac{k}{2})}^{-2\chi-d}}{({\bf q-\frac{k}{2}})^{2}\times\biggl{(}({\bf q+\frac{k}{2}})^{z}+{\bf|q-\frac{k}{2}|}^{z}\biggl{)}}.$ (52)
Next we find ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from numerator, which are relevant to us. Then the above integral $I$ can be split into three parts say $I_{1}$, $I_{2}$ and $I_{3}$. The first part $I_{1}$, after expanding its denominator binomially, is
$I_{1}$ $=-\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{2q^{z}}\times\frac{k_{i}q_{i}}{q^{4}}$
$=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (53)
Next,
$I_{2}$ $=\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}\times\frac{k_{i}q_{i}}{q^{2}}$
$=\frac{\lambda\lambda_{1}D}{2\Gamma}(2\chi+d)\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (54)
In evaluating $I_{1}$ and $I_{2}$ we used relation (42). Finally,
$I_{3}$ $=\frac{2\lambda\lambda_{1}D}{2\Gamma}k^{2}\int_{{\bf q}}\frac{q^{4}q^{-2\chi-d}}{q^{2}\times 2q^{z}}$
$=\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k_{d}k^{2-\chi}}{\chi}.$ (55)
Adding above three contributions we get
$-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}\biggl{[}\frac{2-2\chi-2d}{2}\biggl{]}.$ (56)
Adding Eqs. (49) and (56) we get the total ${\cal O}(\lambda\lambda_{1})$ contribution to self energy at zero-frequency
$\frac{\lambda^{2}D}{\Gamma}\gamma\frac{k_{d}}{2d}\frac{k^{2-\chi}}{\chi}\biggl{[}\frac{4d+2d^{2}+2\chi d+10\chi-6}{d+2}\biggl{]}.$ (57)
Here, $\gamma=\frac{\lambda_{1}}{\lambda}$, as originally defined in the main text. Now adding Eqs. (34), (43) and (57) we get total one-loop contribution to the zero frequency self energy. Now using the definition given in Eq. (29), we obtain the relation
$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{\chi\times d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (58)
Figure 4: One-loop Feynman diagrams that contribute to the zero frequency correlation function.
Next we calculate the one-loop contributions to the correlation function; see Fig. 4.
Figure 4(a) gives the one-loop correction at ${\cal O}(\lambda^{2})$ to the correlation function with a symmetry factor of $2$. Evaluating it at zero frequency, we get
$2\times\Bigl{(}\frac{\lambda}{2}\Bigl{)}^{2}\frac{1}{2!}\times 2\times\frac{1}{\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}C({\bf q},\Omega)\times C({\bf k-q},-\Omega)$
$=\frac{\lambda^{2}}{2\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}4D^{2}\Gamma^{2}q^{-2\chi-d+z}\times\frac{|{\bf k-q}|^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(\Omega^{2}+\Gamma^{2}({\bf k-q})^{2z})}$ (59)
$=\frac{2D^{2}\lambda^{2}}{k^{2z}}\int_{{\bf q},\Omega}\frac{q^{4}\times q^{-4\chi-2d+2z}}{(\Omega^{2}+\Gamma^{2}q^{2z})^{2}}$ (60)
After performing the $\Omega$ integral, we find
[…]
$=\frac{\lambda^{2}D^{2}}{2\Gamma^{3}}k_{d}\frac{1}{(d+3\chi-2)}k^{-d-\chi-2}.$ (62)
Next, Fig. 4(b) is the one-loop correction at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of $2$. Evaluating it at zero frequency, we get
$-2\times\frac{\lambda_{1}^{2}}{2!}\times 2\frac{1}{\Gamma^{2}k^{2z}}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{4}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}q_{m}(q-k)_{n}\times\frac{4D^{2}\Gamma^{2}q^{-4\chi-2d+2z}}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}.$ (63)
Using, $\int_{\Omega}\frac{1}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}=\frac{1}{4\Gamma^{3}q^{3z}}$ above integral reduces to,
$\frac{8\lambda_{1}^{2}D^{2}\Gamma^{2}}{4\Gamma^{5}k^{2z}}\times\frac{k_{i}k_{j}k_{m}k_{n}}{k^{4}}\int_{{\bf q}}q_{i}q_{j}q_{m}q_{n}q^{-4\chi-2d-z}.$ (64)
By using $\chi+z=2$ and identity (40), we get the one-loop contribution to the correlation function at ${\cal O}(\lambda_{1})^{2}$:
$\frac{6\lambda^{2}D^{2}}{\Gamma^{3}}\gamma^{2}\frac{k_{d}}{d(d+2)}\frac{1}{(d+3\chi-2)}k^{-d-\chi-2}.$ (65)
Finally, Fig. 4(c) gives the one-loop contribution to zero frequency correlation function at ${\cal O}(\lambda_{1}\lambda)$ with a symmetry factor of $2$. Evaluating, we find
$-2\times\frac{\lambda\lambda_{1}}{2!}\times 2\frac{1}{\Gamma^{2}k^{2z}}\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}\bigl{[}{\bf q\cdot(q-k)}\bigl{]}\times\frac{4D^{2}\Gamma^{2}q^{-4\chi-2d+2z}}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}$ (66)
$=\frac{8\lambda\lambda_{1}D^{2}\Gamma^{2}}{4\Gamma^{5}k^{2z}}\times\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q}}q_{i}q_{j}q^{2-4\chi-2d-z}$ (67)
$=\frac{2\lambda_{1}\lambda D^{2}}{\Gamma^{3}k^{2z}}\frac{k_{d}}{d}\int_{q}q^{4-4\chi-2d-z+d-1}.$ (68)
In evaluation of the above integral we used relation $\chi+z=2$ and identity (42). After some simplifications, we finally obtain
$\frac{2\lambda^{2}D^{2}}{\Gamma^{3}}\gamma\frac{k_{d}}{d}\frac{k^{-d-\chi-2}}{(d+3\chi-2)}.$ (69)
Now, adding Eqs. (62), (65) and (69), we obtain the total contribution to zero frequency correlation function from all combination of vertices at one loop order, which reads
$\frac{\lambda^{2}D^{2}}{\Gamma^{3}}\frac{k_{d}}{(d+3\chi-2)}k^{-d-\chi-2}\biggl{[}\frac{1}{2}+\frac{2\gamma}{d}+\frac{6\gamma^{2}}{d(d+2)}\biggl{]}.$ (70)
Then equating Eq. (70) with the definition of the zero frequency correlation function given by Eq. (31) we obtain
$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{4d(d+2)(d+3\chi-2)}\biggl{[}d(d+2)+4\gamma(d+2)+12\gamma^{2}\biggl{]}.$ (71)
Comparing the RHS of Eqs. (58) and (71) we obtain a quadratic equation of the roughness exponent $\chi$ as a function of dimension $d$ and dimensionless ratio $\gamma$ which is
$A\chi^{2}+B\chi+C=0.$ (72)
Where,
$A=36\gamma^{2}+(6d+30)\gamma+(3d+6)$
$B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)}$
$C=(d-2)\times\gamma\times\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]},$
as given in the main text.

### v3

@@
 Abstract
-We study a generalized Kardar-Parisi-Zhang (KPZ) equation [D.
-Jana et al, Phys.
+We study a generalized Kardar-Parisi-Zhang (KPZ) equation [Jana et al., Phys.
 Rev.
-E 109, L032104 (2024)], that sets the paradigm for universality in roughening of growing nonequilibrium surfaces without any conservation laws, but with competing local and nonlocal nonlinear effects.
-We show that such a generalized KPZ equation in two dimensions can describe a strong coupling rough or a crumpled surface, in addition to a weak coupling phase.
-The conformation fluctuations of such a rough surface are given by nonuniversal exponents, with orientational long-ranged order and positional short-ranged order, whereas the crumpled phase has positional and orientational short range order.
-Experimental and theoretical implications of these results are discussed.
+E 109, L032104 (2024)] that sets the paradigm for universality in roughening of growing nonequilibrium surfaces without any conservation laws but with competing local and nonlocal nonlinear effects.
+This equation in two dimensions exhibits two distinct strong coupling regimes: a rough phase and a crumpled phase, in addition to a weak coupling phase.
+The conformation fluctuations of such a rough surface are given by nonuniversal scaling exponents, with orientational long-range order and positional short-range order, whereas the crumpled phase has positional and orientational short-range order.
 The Kardar-Parisi-Zhang (KPZ) equation Kardar et al.
@@
 It undergoes a roughening transition between a smooth phase, statistically identical to a surface described by the linear Edwards-Wilkinson (EW) equation Edwards and Wilkinson (1982), and a perturbatively inaccessible strong coupling phase Barabási and Stanley (1995), in dimension $d>2$.
-Perturbative dynamic renormalization group (RG) Barabási and Stanley (1995); Täuber (2014) analysis has been successful in describing this roughening transition in an $\epsilon$-expansion, where $d=2+\epsilon$, with two dimensions (2D) being the lower critical dimension of the model.
+Perturbative dynamic renormalization group (RG) Barabási and Stanley (1995); Täuber (2014) analysis has been successful in describing this roughening transition in an $\epsilon$ expansion, where $d=2+\epsilon$, with two dimensions (2D) being the lower critical dimension of the model.
 In $d\leq 2$, there exists only the strong coupling phase.
@@
 in the scaling limit, where $r=|{\bf x}|$ and $\phi$ is a dimensionless scaling function of its argument Barabási and Stanley (1995).
-In the smooth phase, $z=2$ and $\chi=1-d/2$ Barabási and Stanley (1995) and at the roughening transition at $d=2+\epsilon,\,\epsilon>0$, $z=2$ and $\chi=0$ Barabási and Stanley (1995); Täuber (2014); Janssen et al.
-(1999).
-Failure of RG to capture the strong coupling phase has led to the development of alternative techniques.
+In the smooth phase, $z=2$ and $\chi=1-d/2$ Barabási and Stanley (1995), and at the roughening transition at $d=2+\epsilon,\,\epsilon>0$, $z=2$ and $\chi=0$ Barabási and Stanley (1995); Täuber (2014); Janssen et al.
+(1999), which are accessible within perturbative RG calculations.
+In contrast, the strong coupling rough phase cannot be accessed within standard RG calculations Täuber (2014).
+Failure of RG to capture the rough phase has led to the development of alternative techniques.
 Notable are the self-consistent mode-coupling theories (MCT) Bouchaud and Cates (1993); Bhattacharjee (1998), which have predicted a variety of results on scaling within specific different calculational approaches.
 For example Ref.
-Hwa and Frey (1991) explored the scenario with $d=1$, yielding $z=3/2$, a value known exactly due to the Galilean invariance and a Fluctuation-Dissipation-Theorem Barabási and Stanley (1995).
+Hwa and Frey (1991) explored scaling at $d=1$, yielding $z=3/2$, a value known exactly due to the Galilean invariance and a fluctuation-dissipation theorem Barabási and Stanley (1995).
 Furthermore, Ref.
 Doherty et al.
-(1994) shows that $z$ rises from $z=3/2$ at $d=1$ to 2 around $d_{U}=3.6$, the upper critical dimension of the KPZ equation.
+(1994) showed that $z$ rises from $z=3/2$ at $d=1$ to 222 around $d_{U}=3.6$, where $d_{U}$ is the upper critical dimension of the KPZ equation defined by the dimension in which $z$ touches the value 2 from a smaller numerical value at lower $d<d_{U}$.
 In contrast, Ref.
 Tu (1994) suggests $z<2$ at any finite dimension indicating $d_{U}=\infty$.
-Furthermore, Refs.
-Bhattacharjee (1998); Bhattacharjee and Bhattacharyya (2007); Basu and Frey (2004, 2009a) used MCT to predict $\chi=(4-d)/6,\,z=(8+d)/6$ that tend to agree reasonably with the numerical results Gomes-Filho et al.
+Furthermore, confirming earlier functional RG results of Ref.
+Halpin-Healy (1990), Refs.
+Bhattacharjee (1998); Bhattacharjee and Bhattacharyya (2007); Basu and Frey (2004, 2009) used MCT to arrive, quite separately, at the same dimension-dependent exponents $\chi=(4-d)/6,z=(8+d)/6$, reasonably close to the more recent nonperturbative RG predictions Canet et al.
+(2010).
+This common functional RG and MCT finding suggests explicitly that $d_{U}=4$, supported independently by Ref.
+Fogedby (2005), using a third distinct technique.
+See also Ref.
+Colaiori and Moore (2001), which gives $d_{U}=4$ and the dynamic exponent $z$ in different dimensions.
+Furthermore, Ref.
+Canet and Moore (2007) illustrated the MCT solution for the KPZ equation at $d\geq 2$ that continues to $d_{U}=4$.
+More recent extensive numerical studies, however, appear to differ from the MCT predictions on the scaling exponents Kelling et al.
+(2017); Halpin-Healy (2013); Pagnani and Parisi (2015) and suggest $d_{U}>4$ Schwartz and Perlsman (2012); Pagnani and Parisi (2013); Kim and Kim (2013); Alves et al.
 (2014).
 See also Ref.
-Colaiori and Moore (2001), which gives $d_{c}=4$ and dynamic exponent $z$ in different dimensions.
-More recently, Ref.
-Canet and Moore (2007) revisited the MCT for the KPZ equation and showed that solutions for the MCT equations for the KPZ equation actually have two branches or two distinct universality classes below $d<2$.
-At $d\geq 2$, there is only one branch that continues to $d_{U}=4$, giving scaling exponents closed to those known from numerical studies.
-The dynamics of a KPZ surface depends locally on surface fluctuations, and hence cannot model nonequilibrium surfaces with nonlocal dynamics.
-Nonlocal effects, however, can be important in wide-ranging systems, including biological growth process Santalla and Ferreira (2018), fast nonlocal transport Aharonov and Rothman (1996) and nonlocal stabiliztion of surfaces Krug and Meakin (1991); Nicoli et al.
+Halpin-Healy (2025) for a very recent study on the scaling in the KPZ equation at $d>2$ by combining a variety of techniques.
+Furthermore, the results in Ref.
+Halpin-Healy and Takeuchi (2015) tend to indicate the absence of a finite $d_{U}$.
+The dynamics of a KPZ surface depends locally on surface fluctuations and hence cannot model nonequilibrium surfaces with nonlocal dynamics.
+Nonlocal effects, however, can be important in wide-ranging systems, including biological growth processes Santalla and Ferreira (2018), fast nonlocal transport Aharonov and Rothman (1996), and nonlocal stabilization of surfaces Krug and Meakin (1991); Nicoli et al.
 (2009).
-Recently, a generalized or “active” KPZ equation (hereafter a-KPZ equation) with nonlocal effects as a conceptual model with competing local and nonlocal nonlinear effects has been proposed Jana et al.
-(2024)
-$\frac{\partial h}{\partial t}$ $=$ $\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}$ (2)
-$+$ $\lambda_{1}\int d^{d}rQ_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta.$
+Recently, a generalized KPZ equation (hereafter G-KPZ equation) with nonlocal effects as a conceptual model with competing local and nonlocal nonlinear effects has been proposed Jana et al.
+(2024):
+$\frac{\partial h}{\partial t}$ $=\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}$
+$+$ $\quad\lambda_{1}\int d^{d}r^{\prime}Q_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta.$ (2)
 Here, $Q_{ij}({\bf r})$ denotes the longitudinal projection operator, which in Fourier space is expressed as $Q_{ij}({\bf k})=k_{i}k_{j}/k^{2}$, with ${\bf k}$ being a Fourier wave vector.
 Thus $Q_{ij}({\bf r})$, which is the inverse Fourier transform of $Q_{ij}({\bf k})$, is nonlocal in space.
-Physically, the term $\lambda_{1}\int d^{d}rQ_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))$ contributes to the surface velocity normal to the base plane, $v_{p}=\partial h/\partial t$, and is nonlocal in ${\bm{\nabla}}h$.
+Physically, the term $\lambda_{1}\int d^{d}r^{\prime}Q_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))$ contributes to the surface velocity normal to the base plane, $v_{p}=\partial h/\partial t$, and is nonlocal in ${\bm{\nabla}}h$.
 Furthermore, $\eta({\bf x},t)$ is a zero-mean Gaussian-distributed white noise with a variance $\langle\eta({\bf x},t)\eta({\bf 0},0)\rangle=2D\delta^{d}({\bf x})\delta(t).$ If $\lambda_{1}=0$, Eq.
 (2) reduces to the usual KPZ equation.
-As a consequence of the competition between the local and nonlocal nonlinearities, a-KPZ equation (2) shows macroscopic properties dramatically different from the KPZ equation.
-It can have a stable, sub- or super-logarithmically rough surface in 2D, characterized by nonuniversal scaling exponents.
+Equation (2) serves as a minimal nonequilibrium model to study competition between local and nonlocal nonlinearities.
+For instance, as a consequence of these competitions, the G-KPZ equation (2) shows macroscopic properties dramatically different from the KPZ equation.
+It can have a stable, weak coupling sub- or superlogarithmically rough surface in 2D, characterized by nonuniversal scaling exponents.
 More strikingly and unlike the 2D KPZ equation, it also admits a roughening transition even in 2D, again characterized by nonuniversal exponents.
-At $d>2$, it shows a roughening transition different from that in the KPZ equation Jana et al.
-(2024).
-Equation (2) admits a pseudo-Galilean invariance under the transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}=t$, together with $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, which ensures that the combination $\lambda+2\lambda_{1}$ does not renormalize; see also Ref.
+Furthermore, at $d>2$ it shows a roughening transition different from that in the KPZ equation; see Ref.
+Jana et al.
+(2024), which principally focused on the weak coupling phase behaviors.
+Equation (2) admits a pseudo-Galilean invariance under the transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}=t$, together with $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, which ensures that the combination $\lambda+2\lambda_{1}$ does not renormalize; see Supplemental Material sup , see also Ref.
 Jana et al.
 (2024).
-This further means $\chi+z=2$ exactly so long as nonlinearities remain relevant (in the scaling sense).
-In this Letter, we focus on the strong coupling phases of (2).
-We show that these strong coupling phases can be of two distinct kinds: (i) rough phase with positional short range order (SRO), but orientational long range order (LRO), or (ii) crumpled with both positional and orientational SRO.
-We extract the scaling exponents of the rough phase in $d\geq 2$, which show their nonuniversal nature: The variance $\Delta\equiv\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle\sim L^{2\chi(\gamma)}$, $10<\chi(\gamma)<$ 1 where $L$ is the linear size of the surface and $\overline{h}(t)$ is the average height at time $t$.
-Furthermore, the time-scale of relaxation $\tau(L)\sim L^{z(\gamma)}$.
+This further means $\chi+z=2$ exactly, as long as nonlinearities remain relevant (in the scaling sense).
+In this Letter we focus on the strong coupling phases of (2).
+We show that these strong coupling phases can be of two distinct kinds: (i) a rough phase with positional short-range order (SRO) but orientational long-range order (LRO), or (ii) crumpled with both positional and orientational SRO.
+We extract the scaling exponents of the rough phase in $d\geq 2$.
+The variance $\Delta\equiv\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle\sim L^{2\chi(\gamma)}$, $0<\chi(\gamma)<1$, where $L$ is the linear size of the surface and $\overline{h}(t)$ is the average height at time $t$.
+Furthermore, the timescale of relaxation $\tau(L)\sim L^{z(\gamma)}$.
 Both $\chi$ and $z$ are parametrized continuously by $\gamma\equiv\lambda_{1}/\lambda$, reflecting model parameter-dependent, nonuniversal scaling.
-In the crumpled phase, $\chi>1$.
+In the rough (crumpled) phase, $0<\chi<1\,(\chi>1)$.
+Figure 1: (a) Phases and the RG flow lines in the $g$-$g_{2}$ plane at 2D.
+Stable and unstable regions are divided by the separatrices (red lines).
+Stable (unstable) regions are indicated by flow lines moving towards (away from) the origin.
+The arrows indicate the flow directions, delineating the stable and unstable regions.
+In this flow diagram (0,0) is the only fixed point, which is unstable along the $g$ direction but stable along the $g_{2}$ direction.
+(b) Phases in the $\lambda$-$\lambda_{1}$ plane.
+The green regions are the weak coupling logarithmically rough phase, the white regions are the algebraically rough phase, and the purple regions are the crumpled phase.
+In (a) and (b) the solid red line represents the locus of unstable fixed points where the roughening transition occurs.
+(c) Variation of $\chi$ in 2D in the phase space region where it is positive, with $\gamma$ as obtained from MCT.
+The central region where $\chi>1$ corresponds to a crumpled surface.
+See text.
 We now derive the above results.
@@
 Dimensional analysis allows us to define two effective dimensionless coupling constants:
-$g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d},\,g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}.$ (3)
-Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, and $S_{d}$ is the surface area of a $d$-dimensional sphere.
-Considering $\lambda>0$ without any loss of generality, we notice that from (3) above that $g_{2}$ can be both positive and negative, depending upon the sign of $\gamma$.
+$g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d},\,g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sgn}(\gamma)\sqrt{g_{1}}.$ (3)
+Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, and $S_{d}$ is the surface area of a $d$-dimensional unit sphere.
+Without any loss of generality, we are considering $\lambda>0$.
+We notice that from (3) above that $g_{2}$ can be both positive and negative, depending upon the sign of $\gamma$.
 There are no one-loop corrections to $\lambda,\lambda_{1}$.
@@
 (1977); Jana et al.
-(2024), which allow us to calculate the one-loop RG flow equations for $g$, $g_{1}$ and hence $g_{2}$,
+(2024), which give the one-loop RG flow equations for $g$ and $g_{2}$:
 $\frac{dg}{dl}$ $=$ $g[2-d-g\tilde{A}(\gamma)],$ (4)
@@
 Its stability properties are intriguing.
-It is stable (i.e., attractive) along the $g_{2}$-axis, but unstable (i.e., repulsive) along the $g$-axis.
+It is stable (i.e., attractive) along the $g_{2}$ axis but unstable (i.e., repulsive) along the $g$ axis.
 This indicates the existence of a separatrix, an invariant manifold under RG in the $g$-$g_{2}$ plane, that separates the stable phase from instability.
@@
 Stability requires $dg/dl<0,dg_{2}/dl<0$, so that both $g(l),\,g_{2}(l)$ flow to zero in the long wavelength limit.
-Else when $dg/dl>0,dg_{2}/dl>0$, both $g(l),\,g_{2}(l)$ flow away from (0,0) giving instability.
+Otherwise, when $dg/dl>0,dg_{2}/dl>0$, both $g(l),\,g_{2}(l)$ flow away from (0,0), giving instability.
 These two behaviors are demarcated by the separatrices $\tilde{A}(\gamma)=0$, giving
 $g_{2}=(0.161)\sqrt{g},\;g_{2}=-1.383\sqrt{g}$ (6)
-for $\lambda_{1}>0$ and $\lambda_{1}<0$, respectively, as the separatrices, in $g$-$g_{2}$ plane passing through (0,0).
+for $\lambda_{1}>0$ and $\lambda_{1}<0$, respectively, as the separatrices, in a $g$-$g_{2}$ plane passing through (0,0).
 The RG flow diagram in 2D is shown in Fig.
@@
 The red lines in Fig.
-1(a) are the separatrices given in (6): the RG flow lines in the green regions bordered by the separatrices flow to the origin, giving the “weak coupling phase”.
-Although (0,0) is the stable fixed point in this region, the flow to (0,0) is so slow that $\nu$ and $D$ are infinitely renormalized, giving being super- or sub-logarithmically rough phases, as explored in Ref.
+1(a) are the separatrices given in (6).
+The RG flow lines in the green regions demarcated by the separatrices flow to the origin, giving the “weak coupling phase.” Although (0,0) is the stable fixed point in this region, the flow to (0,0) is so slow that $\nu$ and $D$ are infinitely renormalized, giving super- or sublogarithmically rough phases, as explored in Ref.
 Jana et al.
@@
 As soon as $g(l),\,g_{2}(l)\sim{\cal O}(1)$, which happens for a surface of finite size, the RG flow equations are outside the validity of the perturbative expansions.
-This is the perturbatively inaccessible “strong coupling phase”.
+This is the perturbatively inaccessible “strong coupling phase.”
 We first argue that the region confined between the two separatrices [red lines in Fig.
-1(a)] actually consists of two subregions with distinct scaling properties (instead of just one region with a rough phase, as originally speculated in Ref.
+1(a)] actually consists of two subregions with distinct scaling properties (instead of just one strong coupling phase, as originally speculated in Ref.
 Jana et al.
 (2024)).
-To investigate that let us first consider the bare perturbation theory for $\nu$ in 2D, which is same as the RG calculations for $\nu$ except that now we extend the integrals over wave vectors down to an infrared cutoff $q_{\text{min}}=2\pi/L$.
-Setting $d=2$, we get for the effective $\nu_{e}$
+To investigate that, first consider the bare perturbation theory for $\nu$ in 2D, which is same as the RG calculations for $\nu$ except that now we extend the integrals over wave vectors down to an infrared cutoff $q_{\text{min}}=2\pi/L$.
+Setting $d=2$, we get for the effective $\nu_{e}$,
 $\nu_{e}\approx\nu_{0}+g\nu_{0}\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}\ln(L/a).$ (7)
-We thus find that for $\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}<0$ or $\Bigl{(}\frac{(g_{2})^{2}}{2g}+\frac{5g_{2}}{8\sqrt{g}}\Bigl{)}<0$ i.e.
-$-1.25<\gamma<0$, $\nu_{e}<\nu_{0}$.
-In fact, for sufficiently large $L/a$, $\nu_{e}<0$, giving divergence of $\langle({\bm{\nabla}}h)^{2}\rangle$ implying surface crumpling with positional and orientational SRO.
-In higher dimensions $d>2$, the one loop integrals are finite and hence the lower limit can be extended to 0, corresponding to the thermodynamic limit.
-Then for a range of $\gamma$, $-1.25<\gamma<0$ and low enough bare $\nu_{0}$, $\nu_{e}$ can turn negative, once again implying crumpling.
+We thus find that for $\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}<0$ or $\Bigl{(}\frac{(g_{2})^{2}}{2g}+\frac{5g_{2}}{8\sqrt{g}}\Bigl{)}<0$, $-1.25<\gamma<0$, $\nu_{e}<\nu_{0}$.
+For a crumpled surface, $\langle\delta{\bf n}^{2}\rangle$ diverges for sufficiently large $L/a$, where $\delta{\bf n}$ is deviation of the local normal from a fully flat state.
+In the Monge gauge Nelson et al.
+(1989), to the lowest order in height fluctuations $\delta n\sim{\bm{\nabla}}h$.
+For sufficiently large $L/a$, $\nu_{e}<0$ gives divergence of $\langle({\bm{\nabla}}h)^{2}\rangle$, implying surface crumpling with positional and orientational SRO.
+In higher dimensions $d>2$, the one-loop integrals are finite and hence the lower limit can be extended to 0, corresponding to the thermodynamic limit.
+Then for a range of $\gamma$ given by $-1.25<\gamma<0$ and low enough bare $\nu_{0}$, $\nu_{e}$ can turn negative, once again indicating crumpling.
 We will revisit the region with crumpling below, shown in purple in Fig.
-1(a), by using mode coupling methods, and come to a similar conclusion, giving us confidence about the existence of a crumpled phase.
+1(a), by using mode coupling methods and come to a similar conclusion, giving us confidence about the existence of a crumpled phase.
 See Fig.
 1(b) for a schematic phase diagram in the $\lambda$-$\lambda_{1}$ plane.
-Figure 1: (a) RG flow diagram in the $g$-$g_{2}$ plane in 2D.
-Stable and unstable regions are divided by the separatrices (red lines).
-Stable (Unstable) regions are indicated by flow lines moving towards (away from) the origin.
-The arrows indicate the flow directions, delineating the stable and unstable regions.
-In this flow diagram (0,0) is the only fixed point that is unstable along the $g$-direction, but stable along the $g_{2}$-direction.
-(b) Phases in the $\lambda-\lambda_{1}$ plane.
-The green regions are the weak coupling logarithmically rough phase, the white regions are the algebraically rough phase, and the purple regions are the crumpled phase.
-In (a) and (b) the solid red line represents the locus of unstable fixed points where the roughening transition occurs.
-(c) Variation of $\chi$ in 2D in the phase space region where it is positive, with $\gamma$ as obtained from MCT.
-Central region where $\chi>1$, corresponds to a crumpled surface.
-To study the unstable regions in the RG flow diagrams, we now set up an MCT calculation Bhattacharjee (1998); Basu and Frey (2009b) to extract the scaling behavior; see also Basu (2004); Basu and Bhattacharjee (2005) for applications of MCT in a slightly different context.
-The response $G({\bf k},\omega)$ and correlation $C({\bf k},\omega)$ functions have the following scaling form
-$G({\bf k},\omega)=k^{-z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}=-i\omega+\Sigma(k,\omega),$ (8)
-$C({\bf k},\omega)=k^{-d-2\chi-z}f\biggl{(}\frac{\omega}{k^{z}}\biggr{)}.$ (9)
+To study the unstable regions in the RG flow diagrams, we now set up a one-loop MCT calculation Bhattacharjee (1998); Basu and Frey (2009) and extract the scaling behavior by using it; see also Chattopadhyay et al.
+(2000); Basu (2004); Basu and Bhattacharjee (2005) for applications of MCT in slightly different but related contexts.
+The response $G({\bf k},\omega)$ and correlation $C({\bf k},\omega)$ functions have the following scaling form:
+$G^{-1}({\bf k},\omega)=k^{z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}=-i\omega+\Sigma({\bf k},\omega),$ (8)
+$C({\bf k},\omega)=k^{-d-2\chi-z}f\biggl{(}\frac{\omega}{k^{z}}\biggr{)},$ (9)
 where the self-energy $\Sigma({\bf k},\omega=0)=\Gamma k^{z}$ in the zero frequency limit. Similarly for the correlator in the Lorentzian approximation, we write
@@
 Our argument is that the universal amplitude ratio $\Gamma^{2}/(D\lambda^{2})$ can be written down both from the one-loop diagrammatics of $G^{-1}({\bf k},\omega)$ and $C({\bf k},\omega)$.
-We use $\chi+z=2$, which is an exact result in the strong coupling phases (also consistent with the absence of vertex corrections at the one-loop order).
-Further assuming that $G^{-1}$ and $C$ are dominated by their respective one-loop contributions,which would hold if $z<2$ and $\chi>0$.
+We use $\chi+z=2$, which is an exact result in the strong coupling phases (also consistent with the absence of vertex corrections at the one-loop order), further assuming that $G^{-1}$ and $C$ are dominated by their respective one-loop contributions,which would hold if $z<2$ and $\chi>0$.
 Combining contributions from all the one-loop diagrams for $G({\bf k},\omega=0)$, we obtain sup
 $\frac{\Gamma^{2}}{\lambda^{2}D}$ $=\frac{k_{d}}{\chi\,d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}$
-$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (11)
-Next, considering the one-loop contributions to $C({\bf k},\omega=0)$, we find
+$+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (11)
+From the one-loop contributions to $C({\bf k},\omega=0)$, we find
 $\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}\biggl{[}d(d+2)+4\gamma(d+2)+12\gamma^{2}\biggl{]}}{4d(d+2)(d+3\chi-2)}.$ (12)
-Now comparing RHS of Eqs. (58) and (71), we obtain a quadratic equation of $\chi$
+Now comparing the right-hand side of Eqs. (11) and (12), we obtain a quadratic equation of $\chi$:
 $A\chi^{2}+B\chi+C=0.$ (13)
-Here, $A=36\gamma^{2}+(6d+30)\gamma+(3d+6),\,B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},\,C=(d-2)\gamma\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]}.$ Solving (13)
+Here, $A=36\gamma^{2}+(6d+30)\gamma+(3d+6),\,B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},\,C=(d-2)\gamma\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]}.$ Solving (13),
 $\chi=\frac{-B\pm|\sqrt{B^{2}-4AC}|}{2A}=\chi_{\pm}.$ (14)
-[…]
+We take $\chi=\chi_{+}$, for with $\gamma=0$ only $\chi_{+}$ reduces to $\chi=1/2$ at $d=1$, the known exact value; incidentally, $\chi=\chi_{+}$ reduces to $\chi=1/3$ in 2D and $d_{U}=4$ obtained in an analogous MCT for the pure KPZ equation Bhattacharjee (1998) for $\gamma=0$.
+We now calculate $\chi$ in 2D.
+With $d=2$ in $A$, $B$, and $C$ defined above, we get $C=0$ and $\chi=\frac{-B+|B|}{2A}$, where $B=18\gamma^{2}+22\gamma-4$ and $A=36\gamma^{2}+42\gamma+12$.
+Validity of our MCT requires $\chi>0$.
+This means either $B<0$ and $A>0$, or $B>0$ and $A<0$.
+Notice that $B=0$ gives the separatrix (6) with $B<0$ corresponding to the unstable region in Fig.
+1(a).
+We focus on the region $B<0$, which holds for $-1.383<\gamma<0.161$.
+Within this region,
 $\chi=\frac{-B}{A}=\frac{-18\gamma^{2}-22\gamma+4}{36\gamma^{2}+42\gamma+12}.$ (15)
-On the separatrices (6), $B=0$ and $A=22.77$ corresponding to $\gamma=\gamma_{c1}=0.161$ and $A=19.69$ corresponding to $\gamma=\gamma_{c2}=-1.383$; $A$ changes from these boundary values within the unstable region between the separatrices.
+On the separatrices (6), $B=0$ and $A=22.77$ corresponding to $\gamma=\gamma_{c1}=0.161$ and $A=19.69$, corresponding to $\gamma=\gamma_{c2}=-1.383$; $A$ changes from these boundary values within the unstable region between the separatrices.
 In particular, as $A$, depending upon $\gamma$, decreases, $\chi$ grows, eventually exceeding unity.
-At this point, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface.
-(We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in a RG sense), making our theory given by Eq.
-(2) to breakdown.) Setting $\chi=1$ in (15), we get $\gamma=-1.043,-0.142$.
-Thus in the region $(-1.043<\gamma<-0.142)$ MCT predicts existence of a crumpled phase, which is the purple region in Fig.
-1(a) and is slightly different from the region defined by $\nu_{e}=0$ (see above).
+This gives $\langle({\bm{\nabla}}h)^{2}\rangle$ diverging with $L$, giving a crumpled surface.
+[We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms, e.g., $({\bm{\nabla}}h)^{2n},\,n>1$, not included in (2) should be relevant (in the scaling/RG sense).
+Thus Eq.
+(2), though it predicts the onset of the crumpled phase, cannot be used to study the crumpled phase.] Setting $\chi=1$ in (15), we get $\gamma=-1.043,-0.142$.
+Thus in the region $(-1.043<\gamma<-0.142)$ MCT predicts a crumpled phase, which is the purple region in Fig.
+1(a) and is slightly different from the region defined by the condition $\nu_{e}=0$ [see Eq.
+(7) above].
 Due to the breakdown of our theory, we cannot make any definitive conclusion about any scaling properties in the crumpled region.
 Figure 2: Conjectured “Occam’s razor” global RG flows in the $g$-$g_{2}$ plane.
-Different speculated fixed points characterizing the rough and crumpled phases are marked.
-The broken lines are speculated fixed lines in the rough phase; see text.
+Different speculated fixed points/lines characterizing the rough and crumpled phases are marked.
+The broken lines are speculated, fixed lines in the rough phase; see text.
 […]
-Having ascertained the existence of rough and crumpled phases in 2D, we now speculate on the global fixed point structure including the strong coupling fixed points, which are inaccessible to perturbative RG; see Fig.
+Having ascertained the existence of rough and crumpled phases in 2D, we now speculate on the global fixed point structure, including the strong coupling fixed points, which are inaccessible to perturbative RG; see Fig.
 2.
 The RG flow lines that flow out of the fixed point (0,0) in the unstable region of the phase space should flow to one of these fixed points.
-To analyze these fixed points, we are first guided by the expectation that there should be a strong coupling fixed point at $g\neq 0,g_{2}=0$, which is the expected 2D KPZ string coupling fixed point and should govern the scaling laws for a 2D KPZ surface (which is rough).
-This is marked by a red circle (the other red circle at the origin is another fixed point that is unstable along the $g$-direction, but stable along the $g_{2}$-direction).
+To analyze these fixed points, we are first guided by the expectation that there should be a strong coupling fixed point at $g\neq 0,g_{2}=0$, which is the expected 2D KPZ strong coupling fixed point and should govern the scaling laws for a 2D KPZ surface (which is rough).
+This is marked by a red circle (the other red circle at the origin is another fixed point that is unstable along the $g$ direction but stable along the $g_{2}$ direction).
 In the present model, this fixed point is expected to be unstable against perturbation by a finite $\gamma$ or $g_{2}$.
-On one side of the $g$-axis, there should be a crumpled phase, which we expect to be characterized by a stable crumpled phase fixed point, marked by a blue square in Fig.
+On one side of the $g$ axis, there should be a crumpled phase, which we expect to be characterized by a stable crumpled phase fixed point, marked by a blue square in Fig.
 2.
-On the other side of the $g$-axis, there should be a rough phase.
-Considering the MCT prediction of nonuniversal scaling in the rough phase, we speculate a fixed line marked by broken line in Fig.
+On both sides of the $g$ axis, there should be a rough phase enclosing the $g$ axis.
+Considering the MCT prediction of nonuniversal scaling in the rough phase, we speculate a fixed line in each of these regions marked by broken line in Fig.
 2, corresponding to the rough phase.
-On the side of the $g$-axis with a crumpling fixed, we also expect to have a stable rough phase fixed line for higher negative values of $g_{2}$ (or higher negative values of $\gamma$), giving nonuniversal scaling.
+On the side of the $g$ axis with a crumpling fixed point, we also expect to have another region with a stable rough phase fixed line for higher negative values of $g_{2}$ (or higher negative values of $\gamma$), giving nonuniversal scaling.
 Finally, we generally expect that for larger noises, i.e., with larger values of $D$, for which both $g$ and $g_{2}$ grow, the surface should be less and less ordered.
-This line of reasoning suggests that there should be a crumpling fixed point on the $g$-axis beyond the putative 2D KPZ rough phase fixed point, as well as on the $g_{2}$ axis (marked by a green triangle), indicating eventual crumpling of the surface for high enough $g$ or $g_{2}$.
-These are basically an “Occam”-razor style arguments: Fig.
+This line of reasoning suggests that there should be two roughening fixed points on the $g_{2}$ axis on either side of the $g$ axis (marked by inverted triangles), describing roughening of logarithmically rough crumpling fixed points on the $g$ axis beyond the putative 2D KPZ rough phase fixed point, as well as on the $g_{2}$ axis (marked by green triangles), indicating eventual crumpling of the surface for high enough $g$ or $g_{2}$.
+These are basically “Occam’s razor” style arguments: Fig.
 2 has the simplest topology that naturally reduces to the known RG flow lines for small $g,\,g_{2}$, as shown in Fig.
 1(a).
-At the same time, it gives the putative global flow lines, allowing for a transition to presumed rough and crumpled phases; see Ref.
+At the same time, it gives the putative global flow lines, allowing for a transition to presumed rough and crumpled phases; see Refs.
 Banerjee et al.
 (2019); Mukherjee and Basu (2022a, b) for similar Occam’s razor arguments in different contexts.
-We now consider the model in higher dimensions: $d>2$.
-In $d=2+\epsilon>2$, there are additional fixed points of the RG flows, other than (0,0) (stable fixed point) in the $g$-$g_{2}$ plane.
-In the accessible region i.e.
-$\tilde{A}(\gamma)>0$, (0,0) is the only stable fixed point and the long-wavelength scaling behavior matches that of the Edwards-Wilkinson (EW) equation, characterized by $z=2,\,\chi=(2-d)/2$ Jana et al.
+We now consider the model in higher dimensions $d>2$.
+In $d=2+\epsilon>2$, there are additional fixed points of the RG flows other than (0,0) (stable fixed point) in the $g$-$g_{2}$ plane.
+In the accessible region, i.e., $\tilde{A}(\gamma)>0$, (0,0) is the only stable fixed point and the long-wavelength scaling behavior is identical to that of the Edwards-Wilkinson (EW) equation, characterized by $z=2,\,\chi=(2-d)/2$ Jana et al.
 (2024).
@@
 As $\chi$ exceeds unity, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface.
-(We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in a RG sense), making our theory given by Eq.
-(2) to breakdown.) Setting $\chi=1$, we get a limit on $\gamma$, up to which this roughening transition between a smooth phase and an algebraically rough phase can be observed.
-Beyond this threshold value of $\gamma$ there is a range of $\gamma$ bounded by another threshold such that in the entire range $\tilde{A}(\gamma)$ is small enough to make $\chi>1$, suggesting that the roughening transition in this range of $\gamma$ is actually between a smooth phase and a crumpled phase (and not a rough phase)!
-Overall thus, similar to our discussions for 2D, we expect the strong coupling phase to consist of a rough phase with $10<\chi<$ 1 and a crumpled phase with $\chi>1$.
-Our further theoretical analysis below substantiate this expectation.
-We now briefly discuss the MCT predictions for $d>2$.
+[As before, it suggests that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in the RG sense), making our theory given by Eq.
+(2) to breakdown.] Setting $\chi=1$, we get a limit on $\gamma$, up to which this roughening transition between a smooth phase and an algebraically rough phase can be observed.
+Beyond this threshold value of $\gamma$, there is a range of $\gamma$ bounded by another threshold such that in the entire range $\tilde{A}(\gamma)$ is small enough to make $\chi>1$, suggesting that the roughening transition in this range of $\gamma$ is actually between a smooth phase and a crumpled phase (and not a rough phase)!
+Overall thus, similar to our discussions for 2D, we expect the strong coupling phase to consist of a rough phase with $0<\chi<1$ and a crumpled phase with $\chi>1$.
+See Fig.
+3.
+Figure 3: Schematic phase diagram together with RG flow lines in the $g$-$g_{2}$ plane for $d>2$.
+The blue square at the origin (0,0) is a stable fixed point that controls the scaling of the smooth phase.
+The red circle is an unstable fixed point found in the KPZ equation at $d>2$, which demarcates the strong coupling rough phase and the smooth phase.
+The solid curve lines delineate the strong coupling phases and the smooth phase.
+Within the strong coupling phase regions, the small region bounded between the broken lines, drawn schematically, correspond to the crumpled phase, the rest being the usual rough phase.
+Arrows indicate the RG flow directions.
+We now briefly discuss the MCT predictions for $d>2$ for the strong coupling phases in Fig.
+3.
 Our RG calculations already predicted a smooth phase and a roughening transition presumably to a strong coupling phase.
-The latter however cannot be accessed in the RG calculations.
-Assuming $\chi>0$, we apply MCT to extract the scaling behaviour in the strong coupling phase.
-Focusing on $d=3$, we get $A=36\gamma^{2}+48\gamma+15$, $B=54\gamma^{2}+78\gamma-2.5$ and $C=16\gamma^{2}+24\gamma$, giving $\chi\equiv\chi(\gamma)$.
+The latter, however, cannot be accessed in the RG calculations.
+Assuming $\chi>0$, we apply MCT to extract the scaling behavior in the strong coupling phase.
+Focusing on $d=3$, we get $A=36\gamma^{2}+48\gamma+15$, $B=54\gamma^{2}+78\gamma-2.5$, and $C=16\gamma^{2}+24\gamma$, giving $\chi\equiv\chi(\gamma)$.
 Thus, both $\chi$ and $z=2-\chi$ are nonuniversal, as they are in 2D.
@@
 Similar analysis indicates analogous behavior in the strong coupling phase for any $d>2$.
-We have thus studied the strong coupling phases of a generalized KPZ equation, in which local and nonlocal nonlinear effects compete (in the scaling sense).
-Our results from the RG and bare perturbation theory in 2D suggest that the strong coupling region in the parameter space the strong coupling phase actually is made of two distinct phases, one an algebraically rough phase with $10<\chi<$ 1 corresponding to orientational LRO but positional SRO, and another with $\chi>1$ implying a crumpled phase having both positional and orientational SRO.
+We have thus studied the strong coupling phases of a generalized KPZ equation in which local and nonlocal nonlinear effects compete (in the scaling sense).
+Our results from the RG and bare perturbation theory in 2D suggest that the strong coupling phase region in the parameter space actually is made of two distinct phases, one an algebraically rough phase, with $0<\chi<1$ corresponding to orientational LRO but positional SRO, and another with $\chi>1$, implying a crumpled phase having both positional and orientational SRO.
 These are corroborated by our one-loop MCT calculations.
-While the crumpled phase cannot be systematically explored by our theory, our MCT predicts that the algebraic rough phase is characterized by nonuniversal scaling exponents parametrized by $\gamma$, similar to the nonuniversal exponents in the logarithmically rough phase by the RG calculations.
+The crumpled phase, however, cannot be systematically explored by our theory.
+Our MCT predicts that the algebraic rough phase is characterized by nonuniversal scaling exponents parametrized by $\gamma$, similar to the nonuniversal exponents in the logarithmically rough phase predicted by the RG calculations.
 Similar behavior including a crumpled phase and a rough phase with nonuniversal scaling exponents are predicted by MCT in 3D.
@@
 (2024) showed that chirality has the effect of suppressing the nonlinear instabilities of the RG flow.
-We thus expect that chiral effects together with the nonlocality should be able to suppress the crumpling of the membrane, stabilising orientational order.
-We note that the nonuniversal scaling exponents are a crucial outcome of the lack of vertex renormalization in one-loop MCT.
-We are unable to speculate whether this result protected by any “hidden” symmetry not known at present, or a fortuitous result, or will not hold at higher order perturbation theory.
-Thus, it would be interesting to verify our results numerically, e.g., by using pseudo-spectral methods Basu and Frey (2009a); Basu et al.
-(1998), or by nonperturbative methods Canet et al.
-(2011a, b, 2010).
-Nonetheless, we expect our predictions on the existence of a crumpled phase to hold true.
-Acknowledgement:- A.B.
-thanks the SERB, DST (India) for partial financial support through the TARE scheme [file no.: TAR/2021/000170] and Alexander von Humboldt Stiftung, Germany for partial financial support through the Research Group Linkage Programme (2024).
-Appendix A Renormalization group calculations
-We revisit and reanalyse the renormalization group calculations on the generalized KPZ equation (2) of the main text.
+We thus expect that chiral effects together with the nonlocality should be able to suppress the crumpling of the membrane, stabilizing orientational order.
+We note that the nonuniversal scaling exponents found in the MCT calculations are a crucial outcome of the lack of vertex renormalization in one-loop MCT.
+We are unable to speculate whether this result is protected by any “hidden” symmetry not known at present, or a fortuitous result, or will not hold at higher order perturbation theory.
+Indeed, if there are unequal infinite renormalizations of $\lambda,\,\lambda_{1}$ at some higher loop order, their ratio $\gamma$ will acquire nontrivial fixed point value(s), instead of being marginal.
+In this case, a rough phase would be characterized by isolated fixed point(s), which should replace the fixed lines in the Occam’s razor global RG flow diagram (2).
+Thus, it would be interesting to verify our results numerically, e.g., by using pseudospectral methods Basu and Frey (2009); Basu et al.
+(1998) or by nonperturbative methods Canet et al.
+(2010, 2011a, 2011b).
+However, we expect our predictions on the existence of a crumpled phase to hold true.
+Our theory should apply to nonequilibrium surfaces with other fast relaxing active degrees of freedom living on it.
+We look for future experimental endeavors along this direction.
+Acknowledgement.
+A.B.
+thanks the Alexander von Humboldt Stiftung, Germany for partial financial support through the Research Group Linkage Programme (2024).
+The authors thank T.
+Halpin-Healy for various helpful suggestions.
+Supplemental Material for “Rough or crumpled: Strong coupling phases of a generalized Kardar-Parisi-Zhang surface”
+I Galilean invariance
+We rewrite the generalized KPZ equation Eq. (2) in the main text, which was introduced as a conceptual model incorporating competing local and nonlocal nonlinear interactions,
+$\frac{\partial h}{\partial t}$ $=\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}+\lambda_{1}\int d^{d}r^{\prime}Q_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta,$ (16)
+where $Q_{ij}({\bf x})\equiv\nabla_{i}\nabla_{j}/\nabla^{2}$ is the longitudinal projection operator and takes the form $k_{i}k_{j}/k^{2}$ in the Fourier space; $\bf k$ is a Fourier wavevector.
+Under the transformation $x_{i}^{\prime}\rightarrow x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}\rightarrow t$, where $i$ and $j$ can take values 111 or 222 we obtain,
+$\frac{\partial}{\partial t^{\prime}}\rightarrow\frac{\partial}{\partial t}+(\lambda+2\lambda_{1})\textbf{c}\cdot\bm{\nabla};\quad\bm{\nabla}^{\prime}\rightarrow\bm{\nabla}.$ (17)
+Using these we show below that if the height function $h({\bf x},t)$ transforms as $h^{\prime}({\bf x}^{\prime},t^{\prime})\rightarrow h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, then Eq.
+(16) remains invariant.
+Considering RHS of Eq.
+(16), the first term transforms as
+$\nu\nabla^{\prime 2}h^{\prime}\rightarrow\nu\nabla^{2}h.$ (18)
+The second term transforms as
+$\frac{\lambda}{2}(\bm{\nabla}^{\prime}h^{\prime})^{2}\rightarrow\frac{\lambda}{2}(\bm{\nabla}h)^{2}+\lambda\bm{\nabla}h\cdot\textbf{c}+\frac{\lambda}{2}c^{2}.$ (19)
+The third term transforms as
+$\lambda_{1}\frac{\nabla^{\prime}_{i}\nabla^{\prime}_{j}}{\nabla^{\prime 2}}(\nabla^{\prime}_{i}h^{\prime}\nabla^{\prime}_{j}h^{\prime})\rightarrow\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)+2\lambda_{1}\textbf{c}\cdot\bm{\nabla}h.$ (20)
+Noise term remains invariant. Similarly LHS of Eq. (16) transforms as
+$\frac{\partial h^{\prime}}{\partial t^{\prime}}\rightarrow\frac{\partial h}{\partial t}+(\lambda+2\lambda_{1})\textbf{c}\cdot\bm{\nabla}h+(\lambda+2\lambda_{1})c^{2}.$ (21)
+From the above equations we see that Eq. (16) is invariant.
+II Renormalization group calculations
+We revisit and reanalyze the renormalization group calculations on the generalized KPZ equation (2) of the main text.
 We follow Jana et al.
@@
 where $\hat{h}$ is the dynamic conjugate field and $\mathcal{S}$ is the action functional:
-$S=-\int_{{\bf x},t}\hat{h}D\hat{h}+\int_{{\bf x},t}\hat{h}\bigg{\{}\partial_{t}h-\nu\nabla^{2}h-\frac{\lambda}{2}(\bm{\nabla}h)^{2}-\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)\bigg{\}}.$ (17)
-Using the one loop diagrams from the supplemental material of Ref. Jana et al. (2024) we obtain renormalized $D$ ($D^{<}$) and renormalized $\nu$ ($\nu^{<}$):
+${\mathcal{S}}=-\int_{{\bf x},t}\hat{h}D\hat{h}+\int_{{\bf x},t}\hat{h}\bigg{\{}\partial_{t}h-\nu\nabla^{2}h-\frac{\lambda}{2}(\bm{\nabla}h)^{2}-\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)\bigg{\}}.$ (23)
+Using the one loop diagrams from the supplemental material of Ref. Jana et al. (2024) we obtain renormalized $D$ ($D^{<}$) and renormalized $\nu$ ($\nu^{<}$), we get
 $D^{<}=D\Biggr{[}1+\frac{\lambda^{2}D}{\nu^{3}}k_{d}\biggl{(}\frac{1}{4}+\gamma\frac{1}{d}+\gamma^{2}\frac{3}{d(d+2)}\biggl{)}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{d^{d}q}{q^{2}}\Biggr{]},$ (24)
 $\nu^{<}=\nu\Biggr{[}1+\frac{\lambda^{2}D}{\nu^{3}}k_{d}\biggl{(}\frac{2-d}{4d}+\frac{\gamma}{d(d+2)}\Bigl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\Bigl{)}+\frac{\gamma^{2}}{d}\biggl{)}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{d^{d}q}{q^{2}}\Biggr{]}.$ (25)
-Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, where $S_{d}$ is the surface area of a $d$-dimensional sphere and $\gamma=\frac{\lambda_{1}}{\lambda}$ as defined in the main text.
-Using above equations and Setting $d=2$ in the above integrals, with $b=e^{\delta l}\approx 1+\delta l$ ($l$ is the RG time) and defining dimensionless coupling constant by $g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d}$ as in the main text, we obtain the RG flow equations for $D$ and $\nu$:
+Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, where $S_{d}$ is the surface area of a $d$-dimensional unit sphere and $\gamma=\frac{\lambda_{1}}{\lambda}$ as defined in the main text.
+Using above equations and Setting $d=2$ in the above integrals, with $b=e^{\delta l}\approx 1+\delta l$ ($l$ is the RG time) and defining dimensionless coupling constant by $g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d}$ as in the main text, we obtain the RG flow equations for $D$ and $\nu$
 $\frac{dD}{dl}=D\Biggr{[}z-d-2\chi+g\biggl{(}\frac{1}{4}+\gamma\frac{1}{d}+\gamma^{2}\frac{3}{d(d+2)}\biggl{)}\Biggr{]},$ (26)
 $\frac{d\nu}{dl}=\nu\Biggr{[}z-2+g\biggl{(}\frac{2-d}{4d}+\frac{\gamma}{d(d+2)}\Bigl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\Bigl{)}+\frac{\gamma^{2}}{d}\biggl{)}\Biggr{]}.$ (27)
-Also defining another dimensionless coupling constant, $g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}$ and using above equations we can write the flow equations of $g$ and $g_{2}$:
+Furthermore, defining another dimensionless effective coupling constant, $g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}$ and using above equations we can write the flow equations of $g$ and $g_{2}$:
 $\frac{dg}{dl}=g\Biggr{[}2-d+g\biggl{\{}\frac{1}{4}+\frac{3}{d(d+2)}\gamma^{2}+\frac{\gamma}{d}-\frac{3}{4}\frac{2-d}{d}-\frac{3}{d}\gamma^{2}-\frac{3\gamma}{d(d+2)}\biggl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\biggl{)}\biggl{\}}\Biggr{]},$
@@
 $\frac{dg}{dl}$ $=$ $g[2-d-g\tilde{A}(\gamma)],$ (28)
-$\frac{dg_{2}}{dl}$ $=$ $\frac{g_{2}}{2}[2-d-g\tilde{A}(\gamma)].$ (23)
+$\frac{dg_{2}}{dl}$ $=$ $\frac{g_{2}}{2}[2-d-g\tilde{A}(\gamma)],$ (29)
 where $\tilde{A}(\gamma)=\frac{9}{8}\gamma^{2}+\frac{11}{8}\gamma-\frac{1}{4}$ as given in the main text.
-Appendix B MCT Calculations
-We start with the general definitions of the correlation $C({{\bf k}},\omega)$ and response functions $G({{\bf k}},\omega)$
+III MCT Calculations
+We start with the general definitions of the correlation function $C({{\bf k}},\omega)$ and response function $G({{\bf k}},\omega)$:
 $\delta^{(d)}({\bf k+k^{{}^{\prime}}})\delta(\omega+\omega^{{}^{\prime}})G({{\bf k}},\omega)=\biggl{\langle}\frac{\delta h({\bf k},\omega)}{\delta\eta({\bf k^{{}^{\prime}}},\omega^{{}^{\prime}})}\biggr{\rangle}$ (30)
 $\delta^{(d)}({\bf k+k^{{}^{\prime}}})\delta(\omega+\omega^{{}^{\prime}})C({{\bf k}},\omega)=\langle h({\bf k},\omega)h({\bf k^{{}^{\prime}}},\omega^{{}^{\prime}})\rangle.$ (31)
-The response and correlation functions are assumed to have the following scaling form
+The response and correlation functions are assumed to have the following scaling form in the long time large length scale limit:
 $G({\bf k},\omega)=k^{-z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}$ (32)
 $C({\bf k},\omega)=k^{-d-2\chi-z}f\biggl{(}\frac{\omega}{k^{z}}\biggr{)},$ (33)
-where $z$ is the dynamic exponent and $\chi$ is the roughness exponent. The Dyson’s equation for the self energy in the scaling limit is
+where $z$ is the dynamic exponent and $\chi$ is the roughness exponent; see also main text for their definitions. The Dyson’s equation for the self energy $\Sigma$ in the scaling limit is
 $G^{-1}({\bf k},\omega)=-i\omega+\Sigma({\bf k},\omega).$ (34)
@@
 $\Sigma({\bf k},0)=\Gamma k^{z}.$ (35)
-Here $\Gamma$ is basically renormalised or effective $\nu$. The correlation function in Lorentzian approximation can be written as
+Here $\Gamma>0$ is the effective damping coefficient. The correlation function in Lorentzian approximation can be written as
 $C({\bf k},\omega)=\frac{2D\Gamma k^{-2\chi-d+z}}{\omega^{2}+\Gamma^{2}k^{2z}}.$ (36)
@@
 Figure 4: One-loop Feynman diagrams that contribute to the zero frequency self energy.
-Figure 3(a) shows the one-loop diagrammatic correction to the self energy at ${\cal O}(\lambda)^{2}$ that comes with a symmetry factor of $8$ and contributes
-$\Bigl{(}\frac{\lambda}{2}\Bigl{)}^{2}\times\frac{1}{2!}\times 8\int_{{\bf q},\Omega}{\bf q\cdot(k-q)(q\cdot k)}C({\bf q},\Omega)G({\bf k-q},-\Omega)=\lambda^{2}\int_{{\bf q},\Omega}\frac{{\bf(k-q)\cdot q\times(k\cdot q)}\times 2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(i\Omega+\Gamma({\bf k-q})^{z})}$ (32)
-After performing the $\Omega$ integral and making ${\bf q}\rightarrow{\bf(q+\frac{k}{2})}$ above integral becomes,
-$I$ $=\frac{\lambda^{2}D}{\Gamma}\int_{\bf q}\Bigl{(}{\bf q+\frac{k}{2}}\Bigl{)}^{-2\chi-d+z}{\bf\biggl{[}\Bigl{(}\frac{k}{2}-q\Bigl{)}\cdot\Bigl{(}q+\frac{k}{2}\Bigl{)}\biggl{]}}\frac{{\bf\biggl{[}k\cdot\Bigl{(}q+\frac{k}{2}\Bigl{)}\biggl{]}}}{{\bf\Bigl{(}q+\frac{k}{2}\Bigl{)}}^{z}+{\bf\Bigl{|}q-\frac{k}{2}\Bigl{|}}^{z}}$
-$=\frac{\lambda^{2}D}{2\Gamma}\frac{2\chi+d}{2}\int_{\bf q}\frac{q^{2}({\bf k\cdot q})q^{-2\chi-d}}{q^{z}}\frac{({\bf k\cdot q})}{q^{2}}-\frac{\lambda^{2}D}{2\Gamma}k^{2}\int_{\bf q}\frac{q^{2}}{2q^{z}}q^{-2\chi-d}$ (33)
+Figure 4(a) shows the one-loop diagrammatic correction to the self energy at ${\cal O}(\lambda^{2})$ that comes with a symmetry factor of 888 and contributes
+[…]
+After performing the $\Omega$-integral and making ${\bf q}\rightarrow{({\bf q}+\frac{\bf k}{2})}$, above integral becomes
+$I$ $=\frac{\lambda^{2}D}{\Gamma}\int_{\bf q}\Bigl{(}{\bf q}+\frac{{\bf k}}{2}\Bigl{)}^{-2\chi-d+z}{\biggl{[}\Bigl{(}\frac{\bf k}{2}-{\bf q}\Bigl{)}\cdot\Bigl{|}{\bf q}+\frac{\bf k}{2}\Bigl{|}\biggl{]}}\frac{{\biggl{[}{\bf k}\cdot\Bigl{|}{\bf q}+\frac{\bf k}{2}\Bigl{|}\biggl{]}}}{{\Bigl{|}{\bf q}+\frac{\bf k}{2}\Bigl{|}}^{z}+{\Bigl{|}{\bf q}-\frac{\bf k}{2}\Bigl{|}}^{z}}$
+$=\frac{\lambda^{2}D}{2\Gamma}\frac{2\chi+d}{2}\int_{\bf q}\frac{q^{2}({\bf k\cdot q})q^{-2\chi-d}}{q^{z}}\frac{({\bf k\cdot q})}{q^{2}}-\frac{\lambda^{2}D}{2\Gamma}k^{2}\int_{\bf q}\frac{q^{2}}{2q^{z}}q^{-2\chi-d}.$ (39)
 After simplifications, the first contribution in the second line of the above integral becomes
-$\frac{\lambda^{2}D}{2\Gamma}\times\frac{2\chi+d}{2}\times\frac{k_{d}}{d}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
+$\frac{\lambda^{2}D}{2\Gamma}\,\frac{2\chi+d}{2}\,\frac{k_{d}}{d}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
 Similarly, the second contribution gives
-$\frac{\lambda^{2}D}{2\Gamma}\times\frac{k_{d}}{2}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
+$\frac{\lambda^{2}D}{2\Gamma}\,\frac{k_{d}}{2}\,k^{2}\int_{\bf q}q^{-\chi-1}.$
 Adding the above contributions, we get the ${\cal O}(\lambda^{2})$ correction to the zero frequency self energy
 $I=\frac{\lambda^{2}D}{4\Gamma}\frac{2}{d}k_{d}k^{2-\chi}.$ (40)
-Figure 3(b) shows the one-loop contribution to the self energy at ${\cal O}(\lambda_{1})^{2}$ with a symmetry factor of $8$ and contributes
-$I$ $=\frac{\lambda_{1}^{2}}{2!}\times 8\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}q_{i}(k-q)_{j}\frac{(q-k)_{m}(q-k)_{n}}{(q-k)^{2}}q_{m}k_{n}C({\bf q},\Omega)\times G({\bf k-q},-\Omega)$
-$=\frac{4\lambda_{1}^{2}k_{i}k_{j}k_{n}}{k^{2}}\int_{{\bf q},\Omega}q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}\frac{1}{{\bf(q-k)^{2}}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})}\times\frac{1}{i\Omega+\Gamma{\bf(k-q)}^{z}}$ (35)
-After performing the $\Omega$ integral, we get
+Figure 4(b) shows the one-loop contribution to the self energy at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of 888 and contributes
+[…]
+$=\frac{4\lambda_{1}^{2}k_{i}k_{j}k_{n}}{k^{2}}\int_{{\bf q},\Omega}q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}\frac{1}{{\bf(q-k)^{2}}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})}\times\frac{1}{i\Omega+\Gamma{\bf|k-q|}^{z}}.$ (41)
+After performing the $\Omega$-integral, we get
 $I=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}q^{-2\chi-d}}{{\bf(q-k)}^{2}(q^{z}+{\bf|k-q|}^{z})}.$ (42)
-Then shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$, above integral becomes
-$\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q-\frac{k}{2})_{m}(q-\frac{k}{2})_{n}(q+\frac{k}{2})_{m}}{{\bf(q-\frac{k}{2})}^{2}({\bf|q+\frac{k}{2}|}^{z}+{\bf|q-\frac{k}{2}|}^{z})}{\bf(q+\frac{k}{2})}^{-2\chi-d}.$ (37)
+Then shifting ${\bf q}\rightarrow{\bf q}+\frac{\bf k}{2}$, above integral becomes
+$\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q-\frac{k}{2})_{m}(q-\frac{k}{2})_{n}(q+\frac{k}{2})_{m}}{{({\bf q}-\frac{{\bf k}}{2})}^{2}({|{\bf q}+\frac{\bf k}{2}|}^{z}+{|{\bf q}-\frac{\bf k}{2}|}^{z})}{|{\bf q}+\frac{\bf k}{2}|}^{-2\chi-d}.$ (43)
 Next, we extract the ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from the numerator. Then above integral $I$ can be split into three parts $I_{1}$, $I_{2}$ and $I_{3}$, with
 $I_{1}$ $=-\frac{4\lambda_{1}^{2}D}{2\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}q_{i}q_{j}q_{m}q_{n}q^{2-2\chi-d-z-4}=-\frac{6\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{1}{\chi}k_{d}k^{-\chi+2},$ (44)
-$I_{2}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{j}q_{m}q_{n}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}=\frac{3\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{2\chi+d}{\chi}k_{d}k^{-\chi+2},$ (39)
-where we have used the following identity Yakhot and Orzag (1986),
+$I_{2}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{j}q_{m}q_{n}q^{-2\chi-d}}{2q^{2}q^{z}}\times\frac{2\chi+d}{2}=\frac{3\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{2\chi+d}{\chi}k_{d}k^{-\chi+2},$ (45)
+where we have used the following identity Yakhot and Orszag (1986),
 $k_{i}k_{j}k_{m}k_{n}\int d^{d}qf(q^{2})q_{i}q_{j}q_{m}q_{n}=k_{i}k_{j}k_{m}k_{n}\frac{[\delta_{ij}\delta_{mn}+\delta_{im}\delta_{jn}+\delta_{in}\delta_{jm}]}{d(d+2)}\int d^{d}qf(q^{2})q^{4},$ (46)
@@
 $I_{3}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}k_{i}k_{j}\int_{\bf q}\frac{q_{i}q_{j}q^{2}q^{-2\chi-d}}{4q^{2}q^{z}}=\frac{\lambda_{1}^{2}D}{\Gamma d}\frac{1}{\chi}k_{d}k^{-\chi+2}.$ (47)
-Then, $I_{3}$ is evaluated by using the well-known relation Yakhot and Orzag (1986)
+Then, $I_{3}$ is evaluated by using the well-known relation Yakhot and Orszag (1986)
 $k_{i}k_{j}\int d^{d}qf(q^{2})q_{i}q_{j}=k_{i}k_{j}\times\frac{[\delta_{ij}]}{d}\int d^{d}qf(q^{2})q^{2}.$ (48)
-Adding $I_{1}$, $I_{2}$ and $I_{3}$ we get contribution to zero frequency self energy from vertices $\lambda_{1}$-$\lambda_{1}$
+Adding $I_{1}$, $I_{2}$ and $I_{3}$ we get contribution to zero frequency self energy at ${\cal O}(\lambda_{1}^{2})$:
 $-\frac{\lambda_{1}^{2}D}{\Gamma}\frac{k_{d}}{d}\frac{1}{\chi}k^{2-\chi}\biggl{(}\frac{4-4d-6\chi}{d+2}\biggl{)}.$ (49)
-Figure 3(c) gives one of the two one-loop corrections to the self energy at ${\cal O}(\lambda_{1}\lambda)$,with a contribution
+Figure 4(c) gives one of the two one-loop corrections to the self energy at ${\cal O}(\lambda\lambda_{1})$, with a contribution
 $I=\frac{4\lambda\lambda_{1}}{2!}\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}({\bf q\cdot k})\frac{1}{i\Omega+\Gamma{\bf|k-q|}^{z}}\frac{2D\Gamma q^{-2\chi-d+z}}{\Omega^{2}+\Gamma^{2}q^{2z}}.$ (50)
-After performing the $\Omega$ integral and shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$, we get
-$\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}\frac{(q+\frac{k}{2})^{-2\chi-d}}{{\bf(q+\frac{k}{2})}^{z}+{\bf|\frac{k}{2}-q|}^{z}}.$ (45)
+After performing the $\Omega$-integral and shifting ${\bf q}\rightarrow{\bf q}+\frac{\bf k}{2}$, we get
+$\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}\frac{|{\bf q}+\frac{\bf k}{2}|^{-2\chi-d}}{{|{\bf q}+\frac{\bf k}{2}|}^{z}+|\frac{\bf k}{2}-{\bf q}|^{z}}.$ (51)
 $=-\frac{2\lambda\lambda_{1}D}{\Gamma}\times\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}q_{i}q_{j}\bigl{(}q_{m}+\frac{k_{m}}{2}\bigl{)}\frac{q^{-2\chi-d}}{2q^{z}}\biggl{(}1+\frac{{\bf k\cdot q}}{q^{2}}\biggl{)}^{-(\frac{2\chi+d}{2})}$ (52)
@@
 $=\frac{\lambda\lambda_{1}D}{\Gamma}3k^{2}k_{d}\frac{2\chi+d}{2}\frac{1}{d(d+2)}\int_{{\bf q}}q^{4-2\chi-d-z-2+d-1}-\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k^{2}k_{d}}{d}\int_{{\bf q}}q^{2-2\chi-d-z+d-1}.$ (54)
-In evaluation of these integrals we used identity (40) and (42). After some simplifications, the above integral gives
+In evaluation of these integrals we have used identities (46) and (48). After some simplifications, the above integral gives
 $-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{2d}\frac{k^{2-\chi}}{\chi}\biggl{(}\frac{2-6\chi-2d}{d+2}\biggl{)}.$ (55)
@@
 $I=\frac{4\lambda\lambda_{1}}{2!}\int_{{\bf q},\Omega}{\bf(q\cdot(k-q))}\frac{(q-k)_{i}(q-k)_{j}}{{\bf(q-k)}^{2}}q_{i}k_{j}\times\frac{1}{i\Omega+\Gamma{\bf|k-q|}^{z}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{\Omega^{2}+\Gamma^{2}q^{2z}}.$ (56)
-After performing the $\Omega$ integral above integral is,
+After performing the $\Omega$-integral, above integral becomes
 $\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{{\bf(q\cdot(k-q))}(q-k)_{i}(q-k)_{j}q_{i}q^{-2\chi-d}}{{\bf(q-k)}^{2}\times(q^{z}+{\bf|k-q|}^{z})}.$ (57)
-After shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$ above integral reduces to
-$\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}(q+\frac{k}{2})_{m}(\frac{k}{2}-q)_{m}(q-\frac{k}{2})_{i}(q-\frac{k}{2})_{j}(q+\frac{k}{2})_{i}\times\frac{{\bf(q+\frac{k}{2})}^{-2\chi-d}}{({\bf q-\frac{k}{2}})^{2}\times\biggl{(}({\bf q+\frac{k}{2}})^{z}+{\bf|q-\frac{k}{2}|}^{z}\biggl{)}}.$ (52)
-Next we find ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from numerator, which are relevant to us.
+After shifting ${\bf q}\rightarrow{\bf q}+\frac{\bf k}{2}$, above integral reduces to
+$\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}(q+\frac{k}{2})_{m}(\frac{k}{2}-q)_{m}(q-\frac{k}{2})_{i}(q-\frac{k}{2})_{j}(q+\frac{k}{2})_{i}\times\frac{{\bf(q+\frac{k}{2})}^{-2\chi-d}}{{({{\bf q}-\frac{\bf k}{2}})^{2}\biggl{(}|{\bf q}+\frac{\bf k}{2}}|^{z}+|{\bf q}-\frac{\bf k}{2}|^{z}\biggl{)}}.$ (58)
+Next we calculate the ${\cal O}(k)$ and ${\cal O}(k^{0})$ contributions from numerator, which are relevant.
 Then the above integral $I$ can be split into three parts say $I_{1}$, $I_{2}$ and $I_{3}$.
 The first part $I_{1}$, after expanding its denominator binomially, is
-$I_{1}$ $=-\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{2q^{z}}\times\frac{k_{i}q_{i}}{q^{4}}$
-$=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (53)
+$I_{1}$ $=-\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{2q^{z}}\,\frac{k_{i}q_{i}}{q^{4}}=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (59)
 Next,
-$I_{2}$ $=\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}\times\frac{k_{i}q_{i}}{q^{2}}$
-$=\frac{\lambda\lambda_{1}D}{2\Gamma}(2\chi+d)\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (54)
+$I_{2}$ $=\frac{2\lambda\lambda_{1}D}{\Gamma}k_{j}\int_{{\bf q}}\frac{q^{4}q_{j}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}\times\frac{k_{i}q_{i}}{q^{2}}=\frac{\lambda\lambda_{1}D}{2\Gamma}(2\chi+d)\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}.$ (60)
 In evaluating $I_{1}$ and $I_{2}$ we used relation (48). Finally,
-$I_{3}$ $=\frac{2\lambda\lambda_{1}D}{2\Gamma}k^{2}\int_{{\bf q}}\frac{q^{4}q^{-2\chi-d}}{q^{2}\times 2q^{z}}$
-$=\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k_{d}k^{2-\chi}}{\chi}.$ (55)
+$I_{3}$ $=\frac{2\lambda\lambda_{1}D}{2\Gamma}k^{2}\int_{{\bf q}}\frac{q^{4}q^{-2\chi-d}}{2q^{2}q^{z}}=\frac{\lambda\lambda_{1}D}{2\Gamma}\frac{k_{d}k^{2-\chi}}{\chi}.$ (61)
 Adding above three contributions we get
-$-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}\biggl{[}\frac{2-2\chi-2d}{2}\biggl{]}.$ (56)
+${I_{1}+I_{2}+I_{3}}=-\frac{\lambda\lambda_{1}D}{\Gamma}\frac{k_{d}}{d}\frac{k^{2-\chi}}{\chi}\biggl{[}\frac{2-2\chi-2d}{2}\biggl{]}.$ (62)
 Adding Eqs. (55) and (62) we get the total ${\cal O}(\lambda\lambda_{1})$ contribution to self energy at zero-frequency
@@
 (35), we obtain the relation
-$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{\chi\times d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (58)
+$\frac{\Gamma^{2}}{\lambda^{2}D}=\frac{k_{d}}{\chi\,d}\biggl{[}\frac{\chi}{2}+\gamma\biggl{(}\frac{4d+2d^{2}+2\chi d+10\chi-6}{2(d+2)}\biggl{)}+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (64)
 Figure 5: One-loop Feynman diagrams that contribute to the zero frequency correlation function.
 Next we calculate the one-loop contributions to the correlation function; see Fig. 5.
-Figure 4(a) gives the one-loop correction at ${\cal O}(\lambda^{2})$ to the correlation function with a symmetry factor of $2$. Evaluating it at zero frequency, we get
+Figure 5(a) gives the one-loop correction at ${\cal O}(\lambda^{2})$ to the correlation function with a symmetry factor of 222. Evaluating it at zero frequency, we get
 $2\times\Bigl{(}\frac{\lambda}{2}\Bigl{)}^{2}\frac{1}{2!}\times 2\times\frac{1}{\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}C({\bf q},\Omega)\times C({\bf k-q},-\Omega)$
-$=\frac{\lambda^{2}}{2\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}4D^{2}\Gamma^{2}q^{-2\chi-d+z}\times\frac{|{\bf k-q}|^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(\Omega^{2}+\Gamma^{2}({\bf k-q})^{2z})}$ (59)
+$=\frac{\lambda^{2}}{2\Gamma^{2}k^{2z}}\int_{{\bf q},\Omega}\bigl{(}{\bf q}\cdot({\bf k-q})\bigl{)}^{2}4D^{2}\Gamma^{2}q^{-2\chi-d+z}\times\frac{|{\bf k-q}|^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(\Omega^{2}+\Gamma^{2}|{\bf k-q}|^{2z})}$ (65)
 $=\frac{2D^{2}\lambda^{2}}{k^{2z}}\int_{{\bf q},\Omega}\frac{q^{4}\times q^{-4\chi-2d+2z}}{(\Omega^{2}+\Gamma^{2}q^{2z})^{2}}$ (66)
-After performing the $\Omega$ integral, we find
+After performing the $\Omega$-integral, we find
 […]
-$=\frac{\lambda^{2}D^{2}}{2\Gamma^{3}}k_{d}\frac{1}{(d+3\chi-2)}k^{-d-\chi-2}.$ (62)
-Next, Fig. 4(b) is the one-loop correction at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of $2$. Evaluating it at zero frequency, we get
+Next, Fig. 5(b) is the one-loop correction at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of 222. Evaluating it at zero frequency, we get
 $-2\times\frac{\lambda_{1}^{2}}{2!}\times 2\frac{1}{\Gamma^{2}k^{2z}}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{4}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}q_{m}(q-k)_{n}\times\frac{4D^{2}\Gamma^{2}q^{-4\chi-2d+2z}}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}.$ (68)
-Using, $\int_{\Omega}\frac{1}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}=\frac{1}{4\Gamma^{3}q^{3z}}$ above integral reduces to,
+Using, $\int_{\Omega}\frac{1}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}=\frac{1}{4\Gamma^{3}q^{3z}}$, above integral reduces to
 $\frac{8\lambda_{1}^{2}D^{2}\Gamma^{2}}{4\Gamma^{5}k^{2z}}\times\frac{k_{i}k_{j}k_{m}k_{n}}{k^{4}}\int_{{\bf q}}q_{i}q_{j}q_{m}q_{n}q^{-4\chi-2d-z}.$ (69)
-By using $\chi+z=2$ and identity (40), we get the one-loop contribution to the correlation function at ${\cal O}(\lambda_{1})^{2}$:
+By using $\chi+z=2$ and identity (46), we get the one-loop contribution to the correlation function at ${\cal O}(\lambda_{1}^{2})$:
 $\frac{6\lambda^{2}D^{2}}{\Gamma^{3}}\gamma^{2}\frac{k_{d}}{d(d+2)}\frac{1}{(d+3\chi-2)}k^{-d-\chi-2}.$ (70)
-Finally, Fig. 4(c) gives the one-loop contribution to zero frequency correlation function at ${\cal O}(\lambda_{1}\lambda)$ with a symmetry factor of $2$. Evaluating, we find
+Finally, Fig. 5(c) gives the one-loop contribution to zero frequency correlation function at ${\cal O}(\lambda\lambda_{1})$ with a symmetry factor of 222. Evaluating, we find
 $-2\times\frac{\lambda\lambda_{1}}{2!}\times 2\frac{1}{\Gamma^{2}k^{2z}}\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}\bigl{[}{\bf q\cdot(q-k)}\bigl{]}\times\frac{4D^{2}\Gamma^{2}q^{-4\chi-2d+2z}}{\bigl{(}\Omega^{2}+\Gamma^{2}q^{2z}\bigl{)}^{2}}$ (71)
@@
 Comparing the RHS of Eqs. (64) and (76) we obtain a quadratic equation of the roughness exponent $\chi$ as a function of dimension $d$ and dimensionless ratio $\gamma$ which is
-$A\chi^{2}+B\chi+C=0.$ (72)
+$A\chi^{2}+B\chi+C=0,$ (77)
 where,
-$A=36\gamma^{2}+(6d+30)\gamma+(3d+6)$
-$B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)}$
+$A=36\gamma^{2}+(6d+30)\gamma+(3d+6),$
+$B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},$
 $C=(d-2)\times\gamma\times\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]},$